### v2

# Inferentialist Public Announcement Logic: Base-extension Semantics

Abstract
Proof-theoretic semantics, and base-extension semantics in particular, can be seen as a logical realization of inferentialism, in which the meaning of expressions is understood through their use. We present a base-extension semantics for public announcement logic, building on earlier work giving a base-extension semantics for the modal logic $S5$, which in turn builds on earlier such work for $K$, $KT$, $K4$, and $S4$. These analyses rely on a notion of ‘modal relation’ on bases. The main difficulty in extending the existing B-eS for $S5$ to public announcement logic is to account announcements of the form $[\psi]\phi$, which, in this setting, update the modal relations on bases. We provide a detailed analysis of two classical examples, namely the three-player card game and the muddy children puzzle. These examples illustrate how the inferentialist perspective requires fully explicit information about the state of the participating agents.
Keywords: public announcement logic, modal logic, proof-theoretic semantics, base-extension semantics
1 Introduction
Proof-theoretic semantics (P-tS) is the logical realization of inferentialism [3], within which the meaning of expressions is understood through their use. It is an approach to semantics based on proof — as opposed to truth — where ‘proof’ means ‘valid argument’ — as opposed to derivation in a fixed system. This includes both the semantics of proofs (i.e., validity conditions on ‘arguments’; e.g., [5, 13, 14, 18, 19] among much more) and semantics in terms of proofs (i.e., the meaning of logical constants in terms of consequential relationships; e.g., [15, 16, 17, 11] among much more). We call the first proof-theoretic validity (P-tV) and the second base-extension semantics (Be-S). This nomenclature follows from traditions in the literature, but both branches concern validity and make use of base-extensions in their set-up.
Here we are concerned with B-eS, which establishes a concept of validity for formulae that is grounded in atomic inference. That is, in the simplest case, we take a base $\mathscr{B}$ that is a (countable) set of atomic rules written in the form
$\frac{p_{1}\ldots p_{n}}{p}\quad\mbox{or}\quad p_{1},\dots,p_{n}\Rightarrow p$
where $p$ and each $p_{i}$ is an atom (and $n$ may be 00). We then define a notion of validity relative to $\mathscr{B}$ for formulae $\phi$, $\Vdash_{\mathscr{B}}\phi$, by an inductive definition in which the base case is of the form
$\begin{array}[]{rcl}\Vdash_{\mathscr{B}}p&\mbox{iff}&\mbox{$\vdash_{\mathscr{B}}p$}\end{array}$
where $\vdash_{\mathscr{B}}$ denotes provability using the rules of $\mathscr{B}$. The meaning of complex formulae is then defined inductively in a familiar way; for example,
$\begin{array}[]{rcl}\Vdash_{\mathscr{B}}\phi\wedge\psi&\mbox{iff}&\mbox{$\Vdash_{\mathscr{B}}\phi$ and $\Vdash_{\mathscr{B}}\psi$}\end{array}$
See, for example, [15, 16, 17, 12, 10] for more about the basics of B-eS.
Specifically, this paper is concerned with B-eS for public announcement logic (PAL) (see [1, 4]). It builds directly on two previous papers by the present authors. In the first [6], a base-extension semantics for the family of logics $K$, $KT$, $K4$, and $S4$ was presented by adding a modal accessibility relation between bases that is used to define the validity of modal formulae similarly to Kripke semantics. We give a set of conditions for modal relations in order to ensure soundness and completeness for the logics above. We also show that the conditions given in this paper allow for a Euclidean relation on which the corresponding axiom fails. In the second [6], this work was extended to multi-agent $S5$, by adapting the conditions and allowing for multiple modal relations.
The language of PAL, in the absence of common knowledge, is that of classical propositional logic together with epistemic modalities $K_{a}$ and announcements $[\psi]$. The modality $K_{a}$ is an $S5$ modality corresponding to an agent $a$ and a formula of the form $K_{a}\phi$ is read as ‘$a$ knows $\phi$’. Public announcements are truthful, in the sense that the announced formula has to be true at the time of announcing, and public, in the sense that every agent is aware of the announcement taking place. Announcements (and similar epistemic actions) are generally treated dynamically as an update operation that takes us from a model that represents the initial situation to one that a model of the situation after the announcement. An announcement formula $[\psi]\phi$ is read as $\phi$ holds after the announcement of $\psi$.
In the present paper, we develop the analysis of $S5$ to PAL, the modal logic of knowledge, belief, and public communication — and perhaps the most basic dynamic epistemic logic (DEL). This represents, we suggest, a step towards considering the value of P-tS in the setting of logics motivated by, and deployed in, the analysis of knowledge and public communication.
The main difficulty in extending the existing B-eS for $S5$ is to account for PAL’s announcements of the form $[\psi]\phi$. As already mentioned, in [7], we developed a multi-agent $S5$ that defines the validity of epistemic formula (like $K_{a}\phi$) in reference to relations between bases in much the same way as modalities are handled in Kripke semantics, however with further restrictions on our relations (beyond the frame conditions) to ensure that these relations behave appropriately given that we also have the subset relation. Such relations are called modal relations As already mentioned, announcements are usually treated by an update of the model. In Kripke semantics, for example, an announcement of $\phi$ simply causes all $\neg\phi$ worlds to be discarded from the model. In the Be-S case, the set of bases cannot be changed so simply and we have to restrict our updates to the relations between bases. This will require a more sophisticated update semantics given that we have to ensure that the updated relations are still modal relations. As we will see below, this will cause public announcements to no longer be general in the sense that the same announcement can produce different modal relations if it is updated at different bases (more on that later).
In Section 2, we introduce $S5$ modal logic and PAL together with their Kripke semantics. In Section 3 we discuss some classic examples of public announcements: the three-player card game and the muddy children puzzle. In Section 4, we review the B-eS developed for $S5$ in [7]. This set-up forms the starting point for the B-eS of PAL, which we develop in Section 5. In Section 6, we revisit the examples of the B-eS of PAL. These examples reveal how the inferentialist perspective, as formulated in terms of P-tS, expose the need for an explicit representation of all of the required assumptions. Finally, in Section 7, we discuss possible future work on (other) dynamic epistemic logics and the informational content of B-eS for such logics.
2 Background
2.1 $S5$ and its Kripke Semantics
We define a Kripke semantics for multi-agent $S5$ modal logic (an efficient, yet thorough, presentation of $S5$ in the context of epistemic logic may be found in [9]). We also give a corresponding axiomatic proof system. As we interpret the modal operator epistemically, we write $K_{a}$ to denote the $\square_{a}$-operator corresponding to an agent $a$. Accordingly, a formula $K_{a}\phi$ is read as ‘agent $a$ knows that $\phi$’.
A formula $K_{a}\phi$ holds at a world $w$ iff $\phi$ holds at all worlds $v$ such that $R_{a}wv$, where $R_{a}$ is a (reflexive, transitive, and Euclidean) relation between worlds. The natural way of interpreting this relation is as indistinguishability between agents for the agent $a$, that is, agent $a$ at world $w$ cannot tell whether world $w$ or $v$ is the actual world. Therefore, $K_{a}\phi$ holds at a world $w$ iff at all worlds agent $a$ thinks are possible $\phi$ holds.
Definition 1.
For a set of atomic propositions $P$, atomic propositions $p\in P$ and agents $a$, the language for modal logic $S5$ is generated by the following grammar:
$\phi::=p\mid\bot\mid\neg\phi\mid\phi\to\phi\mid K_{a}\phi$
In Kripke semantics validity is defined with reference to specific worlds and models at which the formula is evaluated. We define models in the usual manner.
Definition 2.
A frame is a pair $F=\langle W,R_{A}\rangle$ in which $W$ is a set of possible worlds, and $R_{A}$ a set of binary relations on $W$, one for each agent $a\in A$. A model $M=\langle F,V\rangle$ is a pair of a frame $F$ and a valuation function $V:P\rightarrow\mathcal{P}(W)$ giving for every propositional $p$ the set of worlds in which $p$ is true s.t. $V(p)\subseteq W$ for every $p$. The relation $R$ is called an $S5$-relation if it is an equivalence relation (i.e., reflexive, transitive, and Euclidean). An $S5$-frame is a frame $F=\langle W,R_{A}\rangle$ in which all $R_{a}\in R_{A}$ are $S5$-relations and an $S5$-model is a model $M=\langle F,V\rangle$ in which $F$ is an $S5$-frame.
Given this, we can give the usual truth conditions for formulae at a world. The validity of a formula is then defined by its being true at all worlds in every model. As we are talking about $S5$ modal logic, we restrict this to $S5$-models.
Definition 3.
Let $F=\langle W,R\rangle$ be an $S5$-frame, $M=\langle F,V\rangle$ be an $S5$-model, and $w\in W$ be a world. That a formula $\phi$ is true at $(M,w)$ — denoted $M,w\vDash\phi$ — is defined as follows:
$M,w\vDash p&\mbox{iff}&w\in V(p)\\ M,w\vDash\phi\to\psi&\mbox{iff}&\mbox{if $M,w\vDash\phi$, then $M,w\vDash\psi$}\\ M,w\vDash\neg\phi&\mbox{iff}&\mbox{not $M,w\vDash\phi$}\\ M,w\vDash\bot&\mbox{iff}&\mbox{never}\\ M,w\vDash K_{a}\phi&\mbox{iff}&\mbox{for all $v$ s.t. $R_{a}wv$, $M,v\vDash\phi$}\\ \end{array}$
If a formula $\phi$ is true at all worlds in a model $M$, we say $\phi$ is true in $M$. A formula is valid iff it is true in all models.
This concludes the Kripke semantics for the epistemic logic underlying PAL. We give an axiomatic proof system that is sound and complete with respect to this semantics. Note that the axioms $T$, $4$, and $5$ correspond to the relation being reflexive, transitive, and Euclidean, respectively. Axiomatic systems for other modal logics can be obtained by only using the corresponding axioms for that relation.
Definition 4.
The proof system $\vdash^{S5}$ for the modal logic $S5$, with $a\in A$ is given by the following axioms and rules:
$\phi\to(\psi\to\phi)&\mbox{\emph{(1)}}\\ (\phi\to(\psi\to\chi))\to((\phi\to\psi)\to(\phi\to\chi))&\mbox{\emph{(2)}}\\ (\neg\phi\to\neg\psi)\to(\psi\to\phi)&\mbox{\emph{(3)}}\\ K_{a}(\phi\to\psi)\to(K_{a}\phi\to K_{a}\psi)&\mbox{\emph{(K)}}\\ K_{a}\phi\to\phi&\mbox{\emph{(T)}}\\ K_{a}\phi\to K_{a}K_{a}\phi&\mbox{\emph{(4)}}\\ \neg K_{a}\phi\to K_{a}\neg K_{a}\phi&\mbox{\emph{(5)}}\\ &\\ \mbox{If $\phi$ and $\phi\to\psi$, then $\psi$}&\mbox{Modus Ponens \emph{(MP)}}\\ \mbox{If $\psi$, then $K_{a}\psi$}&\mbox{Necessitation \emph{(NEC)}}\end{array}$
A formula is provable in S5, written $\vdash^{S5}\phi$, if it follows from these axioms and rules. Just the axioms (1)-(3) together with the rule (MP) constitute a proof system for classical logic. Proof systems for the modal logics other than $S5$ are obtained by taking only the corresponding axioms (e.g., modal logic $K4$ results from adding (K), (4), and (NEC) to the classical proof system).
2.2 Public Announcement Logic
To obtain the language for PAL, we enrich the language of multi-agent $S5$ modal logic with announcement operators of the form $[\phi]\psi$, which are read as ‘after a (truthful and public) announcement of $\phi$, $\psi$ holds’. Note that $[\phi]$ is a box-like operator and so it should read ‘after every announcement of $\phi$’; however, these two readings are equivalent as announcements are partial functions (i.e., there is at most one announcement of $\phi$).
Definition 5.
For a set of atomic propositions $P$, atomic propositions $p\in P$ and agents $a$, the language of $PAL$ is generated by the following grammar:
$\phi:=p\mid\bot\mid\neg\phi\mid\phi\to\phi\mid K_{a}\phi\mid[\phi]\phi$
The Kripke semantics of PAL is obtained by enriching that of multi-agent $S5$ with an update operation that restricts the worlds to only those in which the announced formula is true. Announcing a formula is therefore seen as a model-transformation operation that results in a new Kripke model representing the state of knowledge of the agents after the announcement takes places.
Following that idea, we write $M|\phi$ for the model corresponding to $M$ restricted to worlds at which $\phi$ holds.
Definition 6.
For any model $M=\langle F,V\rangle$ with $F=\langle W,R\rangle$ and formula $\phi$, we define $M|\phi=\langle F^{\prime},V^{\prime}\rangle$ with $F^{\prime}=\langle W^{\prime},R^{\prime}\rangle$ in the following way:
$W^{\prime}&=&\{w\mid w\in W\mbox{ and }M,w\vDash\phi\}\\ R^{\prime}&=&R\cap(W^{\prime}\times W^{\prime})\\ V(p)^{\prime}&=&V(p)\cap W^{\prime}\\ \end{array}$
Given this, we can define validity, with a formula of the form $[\phi]\psi$ being true at a world if $\psi$ is true at that world after the announcement of $\phi$. We add this condition to out truth conditions in Definition 3.
Definition 7.
Let $F=\langle W,R\rangle$ be an $S5$-frame, $M=\langle F,V\rangle$ an $S5$-model and $w\in W$ a world. That a formula $\phi$ is true at $(M,w)$—denoted $M,w\vDash\phi$— is defined as follows:
$M,w\vDash p&\mbox{iff}&w\in V(p)\\ M,w\vDash\phi\to\psi&\mbox{iff}&\mbox{if $M,w\vDash\phi$, then $M,w\vDash\psi$}\\ M,w\vDash\neg\phi&\mbox{iff}&\mbox{not $M,w\vDash\phi$}\\ M,w\vDash\bot&\mbox{iff}&\mbox{never}\\ M,w\vDash K_{a}\phi&\mbox{iff}&\mbox{for all $v$ s.t. $R_{a}wv$, $M,v\vDash\phi$}\\ M,w\vDash[\phi]\psi&\mbox{iff}&\mbox{$M,w\vDash\phi$ implies $M|\phi,w\vDash\psi$}\end{array}$
If a formula $\phi$ is true at all worlds in a model $M$, we say $\phi$ is true in $M$. A formula is valid iff it is true in all models.
A sound and complete proof-system for PAL is given by the following axiomatic system:
Definition 8.
We add the following axioms to the axiomatic system for $S5$ modal logic in Definition 4:
$[\phi]p\leftrightarrow\phi\to p&\mbox{atomic permanence}\\ [\phi]\bot\leftrightarrow\phi\to\bot&\mbox{announcement and bot}\\ [\phi](\psi\to\chi)\leftrightarrow[\phi]\psi\to[\phi]\chi&\mbox{announcement and implication}\\ [\phi]K_{a}\psi\leftrightarrow\phi\to K_{a}[\phi]\psi&\mbox{announcement and knowledge}\\ [\phi][\psi]\chi\leftrightarrow[\phi\wedge[\phi]\psi]\chi&\mbox{announcement composition}\\ \mbox{If $\psi$, then $[\phi]\psi$}&\mbox{necessitation of announcements}\end{array}$
A formula is provable in PAL, written $\vdash\phi$, if it follows from these axioms and rules.
3 Examples of Public Announcements
It is useful here to discuss some examples of how the reasoning of PAL can be applied. This shows how the standard semantics for PAL works and will allow us to highlight, when we return to the examples in Section 6, the differences to the base-extension semantics treatment of PAL developed later in this paper.
The examples discussed in this paper and their treatment in the standard semantics is taken from [4].
Our first example will be a simple card game between three agents and our second example, the most famous puzzle concerning PAL, is the muddy-children puzzle. We have chosen these examples as they highlight different aspects of how informational content is encoded in our semantics and how that differs with the treatment of these examples in Kripke semantics, as we will discuss in Section 3.
3.1 Three-player Card Game
Example 1 (Card game).
Three players, Anne, Bob, and Cath, each draw a card from a deck of three cards, 0, 1, and 2. This is known among the agents. Anne draws 0, Bob 1, and Cath 2. Only the players themselves know which card they have drawn. Anne announces ‘I do not have card 1’.
Before it is decided who draws which card, there are six possible distributions of the card draw. We write, for example, 012 for the case in which Anne draws 0, Bob draws 1, and Cath 2 (i.e., the actual case in our example).
Before, however, looking at the thought processes of Anne, Bob, and Cath, we need to analyse the framework in which the game takes place. There are some important rules that must be obeyed in interpretation of the example. Every agent draws exactly one card. By the rules of the game, no agent can draw two cards or no card at all. Conversely, every card is drawn by exactly one agent. There are no cards left out of the deck and there are no multiples of any card. Lastly, every agent can see their own card but not the cards of the other agents.
Given these rules, we can start drawing some conclusions: For every agent, there are certain draws between which they cannot decide. For example, Anne cannot decide between 012 and 021, as they can only say that they themselves have 0 and that Bob and Cath each have one of the other two cards. For the same reasons, Cath cannot decide between 012 and 102. In fact, these are the only draws she considers possible at 012.
Now we go forward to Anne’s announcement ‘I do not have card 1’. This announcement contradicts 102 and so Cath no longer considers it. This leaves 012 as the only possible draw Cath considers and they now know exactly who has which card.
[…]
Figure 1: The initial Kripke model $M$ for the card game example.
Following [4], we use $N_{i}$ for the atomic sentences so that ‘agent $i$ has card $N$’, for example at world $012$, we have $0_{a},1_{b},$ and $2_{c}$. Note that the way our model is set-up guarantees that the rules of the game, as discussed above, hold at all the worlds.
Anne’s announcement of ‘I do not have card 1’ corresponds to $\neg 1_{a}$. So the announcement, as expected, will restrict our model to those worlds at which Anne does not have card $1$. Meaning we delete the worlds $102$ and $120$. Similarly, the relation is restricted to only the remaining worlds. This results in the model $M^{\prime}$ of Figure 2.
[…]
Figure 2: The Kripke model $M^{\prime}$ for the card game example after Anne’s announcement ‘I do not have card 1’.
Given this we can now show that after Anne’s announcement Cath has enough information to identify the draw that took place. We can see that $M^{\prime},012\models K_{c}(0_{a}\wedge 1_{b}\wedge 2_{c})$ and so $M,012\models[\neg 1_{a}]K_{c}(0_{a}\wedge 1_{b}\wedge 2_{c})$.
As we can see, the Anne’s announcement leads to a model in which Cath can figure out who has which card, just as we expected.
3.2 The Muddy Children Puzzle
Again, like with the card game example, we start by informally describing the muddy children puzzle and then show how this puzzle can be solved using Kripke models.
Example 2 (Muddy Children).
A group of children playing outside are called by their father. Some of them have become dirty from playing outside and they have mud on their foreheads. The children cannot see their own forehead but the foreheads of all the other children. The children are, of course, perfect logicians.
The father proclaims: ‘At least one of you has mud on their forehead. Will those with mud on their forehead come forward to get cleaned?’. The children do not like being cleaned and will avoid so if possible (i.e., they do not have muddy foreheads). If no child steps forward the father will repeat this request until the muddy children step forward.
If $m$ of the $n$ children are muddy, the children will step forward after the father has made his request $m$ times.
We are going to focus on the case in which there are three children, two of which are muddy. Let us call the children Anne, Bob, and Cath as in the card game example above. This gives us a specific example to reason through and the general case follows the exact same reasoning.
So how are the children capable of finding out whether they are muddy or not without seeing their own foreheads? Let us take Anne’s point of view (Bob’s, of course, will be analogous). Anne sees that Bob is muddy and that Cath is not muddy. The father now states that at least one of them is muddy. So Anne considers that Bob can also see that Cath is not muddy and so if Anne was not muddy, Bob would realize that they were the only muddy one. When the father repeats his request without Bob having stepped forward, Anne realizes that Bob has to see another muddy person. That can only be Anne herself though and so Anne (and Bob) step forward as they now know that they are the muddy ones.
Similarly to the card game example, we can represent this using Kripke models and by treating the actions of the father and the children as public announcements. The atomic sentences are $m_{a},m_{b}$, and $m_{c}$ for Anne, Bob and Cath are muddy, respectively. Given that we have three atomic sentences we can represent the different worlds in a cube (see Figure 3).
[…]
Figure 3: The initial Kripke model for the muddy children puzzle ahead of the father’s announcements.
We can model the father’s announcement that some of the children are muddy by an announcement of $m_{a}\vee m_{b}\vee m_{c}$. For readability, let $\phi=m_{a}\vee m_{b}\vee m_{c}$. This removes the world in which no child is muddy and all relations between it and the other worlds (see Figure 4).
$m_{b}$$m_{c}$$m_{a}$$m_{a},m_{b}$$m_{a},m_{c}$$m_{b},m_{c}$$m_{a},m_{b},m_{c}$$b$$c$$a$$c$$a$$b$$c$$b$$a$
Figure 4: The Kripke model for the muddy children puzzle after the announcement of $m_{a}\vee m_{b}\vee m_{c}$ (i.e., that some children are muddy).
Now the father repeats his request for the children to get cleaned. Above we treated the children not stepping forward to get cleaned as an announcement that they do not know whether they are muddy or not. This corresponds to the announcement of the formula $\psi=\neg(K_{a}m_{a}\vee K_{a}\neg m_{a})\wedge\neg(K_{b}m_{b}\vee K_{b}\neg m_{b})\wedge\neg(K_{c}m_{c}\vee K_{c}\neg m_{c})$. So, when the children fail to step up after the first request by the father, we update our model of Figure 4 with that announcement. This results in Figure 5.
$m_{a},m_{b}$$m_{a},m_{c}$$m_{b},m_{c}$$m_{a},m_{b},m_{c}$$c$$b$$a$
Figure 5: The Kripke model for the muddy children puzzle after the father has to repeat his request (modelled by the announcement of $\neg(K_{a}m_{a}\vee K_{a}\neg m_{a})\wedge\neg(K_{b}m_{b}\vee K_{b}\neg m_{b})\wedge\neg(K_{c}m_{c}\vee K_{c}\neg m_{c})$)
In the model of Figure 5, we can see that at the world with $m_{a},m_{b}$ Anne and Bob do not have any other worlds they consider and so $K_{a}m_{a}$ and $K_{b}m_{b}$. So from our initial model $[\phi][\psi]K_{a}m_{a}\wedge K_{b}m_{b}$ holds. Cath, of course, still believes it is possible that they themselves are muddy as well, as they still think all of them might be muddy (at least until Anne and Bob step forward).
Note that this reasoning is not depending on the number of total children. Regardless of $n$, this reasoning allows the $m$ actually muddy children to realize that they are muddy in $m$ steps, as every repetition of the father’s request will remove those worlds with the least muddy children until there are no more alternatives to the actual world.
4 Base-extension Semantics for $S5$
We now give a base-extension semantics for multi-agent $S5$, as presented in [7]. This approach builds on that for the systems $K$, $KT$, $K4$, and $S4$, as presented in [6]. The correctness of base extension semantics for $S5$ is established relative to its Kripke semantics [7]. This is the starting point from which we develop the base-extension semantics for PAL.
Definition 9.
For atomic propositions $p$ and agents $a$, the language for the epistemic modal logic is generated by the following grammar:
$\phi:=p\mid\bot\mid\phi\to\phi\mid K_{a}\phi$
Base rules and bases build the foundation of validity, we have to define them before moving to validity conditions.
Definition 10.
A base rule is a pair $(L_{j},p)$, where $L_{j}=\{p_{1},\dots,p_{n}\}$, is a finite (possibly empty) set of atomic formulae and $p$ is also an atomic formula. Generally, a base rule will be written as $p_{1},\dots,p_{n}\Rightarrow p$ ($\Rightarrow p$ for axioms). A base $\mathscr{B}$ is any countable collection of base rules. We call the set of all bases $\Omega$ and $\overline{\mathscr{B}}$ is the closure of the empty set under the rules in $\mathscr{B}$.
A base is called inconsistent if $p\in\overline{\mathscr{B}}$ for all $p$. All other bases are consistent.
Before we define validity, we define the notion of modal relation. We impose four conditions on modal relations. Conditions (a) and (b) separate consistent and inconsistent bases. Condition (c) guarantees that if there is a relation from a base $\mathscr{B}$ to base $\mathscr{C}$, then for every consistent superset base of $\mathscr{B}$, there is a superset base of $\mathscr{C}$ between which the relation holds. Analogously, condition (d) guarantees that if there is a relation from a base $\mathscr{B}$ to base $\mathscr{C}$, then for every subset base of $\mathscr{B}$, there is a subset base of $\mathscr{C}$ between which the relation holds. The conditions are illustrated in Figures 6, 7, and 8. See [7] for a more detailed discussion of these conditions.
Definition 11.
A relation $\mathfrak{R}_{a}$ on the set bases $\Omega$ is called a $S5$-modal relation iff, for all $\mathscr{B}$,
  1. (a)
    if $\mathscr{B}$ is inconsistent, then there is a $\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$ and $\mathscr{C}$ is inconsistent and, for all $\mathscr{D}$, $\mathfrak{R}_{a}\mathscr{B}\mathscr{D}$ implies that $\mathscr{D}$ is also inconsistent.
  2. (b)
    if $\mathscr{B}$ is consistent, then for all $\mathscr{C}$, s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$, $\mathscr{C}$ is also consistent
  3. (c)
    for all $\mathscr{C}$, if $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$, then, for all consistent $\mathscr{D}\supseteq\mathscr{B}$, there is an $\mathscr{E}\supseteq\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{D}\mathscr{E}$
  4. (d)
    for all consistent $\mathscr{C}$, if $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$, then, for all $\mathscr{D}\subseteq\mathscr{B}$, there are $\mathscr{E}\subseteq\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{D}\mathscr{E}$.
A modal relation $\mathfrak{R}$ is called a $S5$-modal relation iff $\mathfrak{R}$ is reflexive, transitive, and Euclidean.
We define validity at a base given a set of model relations. As mentioned in Section 1, validity is defined inductively with the base case for atoms is defined in terms of provability from the rules in the base. The modal relations are used for formulae of the form $K_{a}\phi$ in a manner analogous to Kripke semantics. General validity is defined as being valid at all bases given any modal relation.
Definition 12.
For the epistemic modal logic, we define validity at a base $\mathscr{B}$ given a set of agents $A$ and a set of $S5$-modal relations $\mathfrak{R}_{a}$, one for every agent $a\in A$, denoted by $\mathfrak{R}_{A}$ as follows:
$\Vdash_{\mathscr{B},\mathfrak{R}_{A}}p&\mbox{iff}&\mbox{$p$ is in every set of basic sentences closed under $\mathscr{B}$}\\ &&\mbox{(i.e., iff $p\in\overline{\mathscr{B}}$)}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to\psi&\mbox{iff}&\mbox{$\phi\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}\bot&\mbox{iff}&\mbox{$\Vdash_{\mathscr{B},\mathfrak{R}_{A}}p$ for every basic sentence $p$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\phi&\mbox{iff}&\mbox{for all $\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$, $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$}\\ &&\\ \mbox{For non empty $\Gamma$:}&&\\ \Gamma\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\psi$ for all $\psi\in\Gamma$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$}\\ \end{array}$
Given a set of agents $A$, a formula $\phi$ is valid, written as $\Vdash^{A}\phi$, iff $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ for all modal bases $\mathscr{B}$ and set of modal relations $\mathfrak{R}_{A}$.
[…]
Figure 6: Illustration of Definition 11 (a) on the left and (b) on the right
[…]
Figure 7: Illustration of Definition 11 (c) with dotted lines representing the subset relation
$\mathscr{D}$$\mathscr{B}$$\mathscr{C}$Longrightarrow $\mathscr{D}$$\mathscr{E}$$\mathscr{B}$$\mathscr{C}$
Figure 8: Illustration of Definition 11 (d) with dotted lines representing the subset relation
We prove some lemmas important in establishing the soundness and completeness results, as they will be generalized to the semantics for PAL in Section 5. More details on how these are used in the proof of soundness and completeness can be found in [7].
Lemma 1.
If $\Gamma\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ and $\mathscr{B}\subseteq\mathscr{C}$, then $\Gamma\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$.
Proof.
This proof goes by induction on the complexity of $\phi$. If $\Gamma$ is not empty, this follows immediately from the validity conditions.
For empty $\Gamma$, the propositional cases remain unchanged from the proof for classical propositional logic in [16]. The new case is the case in which $\phi=K_{a}\psi$. In order for $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\psi$, at all $\mathscr{D}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{D}$, $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\psi$. By condition (d) of Definition 11, we know that for every $\mathscr{E}$ s.t. $\mathfrak{R}_{a}\mathscr{C}\mathscr{E}$, there is such a $\mathscr{D}$ with $\mathscr{D}\subseteq\mathscr{E}$. By induction hypothesis $\Vdash_{\mathscr{E},\mathfrak{R}_{A}}\psi$ and, since this holds for all $\mathscr{E}$ s.t. $\mathfrak{R}_{a}\mathscr{C}\mathscr{E}$, we conclude $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}K_{a}\psi$. ∎
Additionally, the conditions of Definition 11 guarantees ex falso quod libet in the sense that any formula $\phi$ follows from $\bot$.
Lemma 2.
For all inconsistent bases $\mathscr{B}$ and $\mathfrak{R}_{A}$, $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$, for all $\phi$.
Proof.
This proof goes by induction on the complexity of the formula $\phi$. For propositional $\phi$ this follows immediately from the validity conditions in Definition 12. For the case in which $\phi=K_{a}\psi$, the condition (a) ensures that any inconsistent base can only have access to other inconsistent bases and, by induction hypothesis, $\psi$ holds at these bases. ∎
We define maximally-consistent bases as they behave interestingly.
Definition 13.
A base $\mathscr{B}$ is maximally-consistent iff it is consistent and for every base rule $\delta$, either $\delta\in\mathscr{B}$ or $\mathscr{B}\cup\{\delta\}$ is inconsistent.
We can show that Kripke-like validity conditions hold for maximally-consistent bases as they do not have consider extensions.
Lemma 3.
For any set of $S5$-modal relations $\mathfrak{R}_{A}$ and maximally-consistent base $\mathscr{B}$, the following hold:
  • $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}\bot$, and
  • $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to\psi$ iff $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ or $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi$.
Proof.
As the validity conditions for $\to$ and $\bot$ do not depend on $\mathfrak{R}_{A}$ the proof remains unchanged from [8]. ∎
Additionally, for any formula $\phi$, at a maximally-consistent base either it or its negation will hold.
Lemma 4.
For every formula $\phi$, maximally-consistent base $\mathscr{B}$, and set of $S5$-modal relation $\mathfrak{R}_{A}$, either $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ or $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to\bot$.
Proof.
Consider the case in which $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$. Any base $\mathscr{C}$ s.t. $\mathscr{C}\supset\mathscr{B}$ is inconsistent and so $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\bot$ and so $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to\bot$. ∎
Lemma 5.
For every formula $\phi$, if there is a base $\mathscr{B}$ s.t. $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$, then there is a maximally-consistent base $\mathscr{C}\supseteq\mathscr{B}$ with $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$.
Proof.
The proof goes by induction on the complexity of the formula $\phi$. The proof for the classical cases follows the proof of Lemma 3.3 and 3.4 in [8].
We consider the case in which $\phi=K_{a}\psi$. Since $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\psi$, there is a $\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$ and $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}\psi$. By our induction hypothesis, we know that there is a maximally-consistent $\mathscr{C}^{*}\supseteq\mathscr{C}$. By symmetry, $\mathfrak{R}_{a}\mathscr{C}\mathscr{B}$ and so, by condition (c) of Defintition 11, there is a $\mathscr{B}^{\prime}\supseteq\mathscr{B}$ s.t. $\mathfrak{R}_{a}\mathscr{C}^{*}\mathscr{B}^{\prime}$. Again by symmetry, we have $\mathfrak{R}_{a}\mathscr{B}^{\prime}\mathscr{C}^{*}$. If $\mathscr{B}^{\prime}$ is maximally-consistent, let $\mathscr{B}^{\prime}$ be our $\mathscr{B}^{*}$. Otherwise, note that, for any maximally-consistent $\mathscr{B}^{*}\supseteq\mathscr{B}^{\prime}$, we have $\mathfrak{R}_{a}\mathscr{B}^{*}\mathscr{C}^{*}$ because of condition (c) again and because $\mathscr{C}^{*}$ is maximally-consistent. Finally, since $\nVdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\psi$, we have $\nVdash_{\mathscr{B}^{*},\mathfrak{R}_{A}}K_{a}\psi$ [^1] It might be of interest to note here that this proof relies on $\mathfrak{R}_{a}$ being symmetric. This suggests that it cannot be used to show that this property holds for arbitrary modal relations. More conditions might be required for this lemma to hold for non-symmetric modal relations.
The following soundness and completeness theorem for multi-agent $S5$’s base-extension semantics relative to its Kripke semantics has been established in [7]:
Theorem 1.
The base-extension semantics for multi-agent modal logic $S5$ is sound and complete with respect to the multi-agent modal logic $S5$
5 Base-extension Semantics for PAL
In order to develop a base-extension for PAL, we must first give a way of representing the model-update of PAL’s Kripke semantics in base-extension semantics. The set of bases, however, is given by the propositional letters and the shape of our base rules. So, we look at updating the modal relations $\mathfrak{R}_{a}$ instead.
Given a relation $\mathfrak{R}_{a}$ and an announcement of a formula $\phi$, we want to create a new relation $\mathfrak{R}_{a}|\phi$ that corresponds to $\mathfrak{R}_{a}$ after the announcement of $\phi$ much like $M|\phi$ corresponds to $M$.
An obvious candidate for such an $\mathfrak{R}_{a}|\phi$ might be that subset of $\mathfrak{R}_{a}$ where for all $\mathscr{B}$, $\mathscr{C}$ s.t. $\mathfrak{R}_{a}|\phi\mathscr{B}\mathscr{C}$, $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ iff $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$. This is, however, not guaranteed to be a modal relation.
For example, take some $\mathscr{B}^{\prime}\subset\mathscr{B}$ and $\mathscr{C}^{\prime}\subset\mathscr{C}$ s.t. $\mathscr{B}^{\prime}\not\subseteq\mathscr{C}^{\prime}$, $\mathfrak{R}_{a}\mathscr{B}^{\prime}\mathscr{C}^{\prime}$ and where $\Vdash_{\mathscr{B}^{\prime},\mathfrak{R}_{A}}\phi$, but $\nVdash_{\mathscr{C}^{\prime},\mathfrak{R}_{A}}\phi$. Let $\mathscr{C}^{\prime}$ be the only base s.t. $\mathfrak{R}_{a}\mathscr{B}^{\prime}\mathscr{C}^{\prime}$ other than $\mathscr{B}^{\prime}$ itself. So we will not have $\mathfrak{R}_{a}|\phi\mathscr{B}^{\prime}\mathscr{C}^{\prime}$. In this case $(d)$ of Definition 11 does not hold and $\mathfrak{R}_{a}|\phi$ is not a modal relation (see Figure 9).
[…]
Figure 9: Updating $\mathfrak{R}_{a}$ (left) by simply disconnecting $\phi$ and non-$\phi$ bases need not result in a $\mathfrak{R}_{a}|\phi$ that is a modal relation (right)!
There is no clear way of forcing a relation like $\mathfrak{R}_{a}|\phi$ to be a modal relation without adding relations between bases that disagree on $\phi$ or by terminating relations between bases that agree on $\phi$. So, our solution to this problem is to consider different $\mathfrak{R}_{a}|\phi$ depending on at which base we are evaluating the announcement. Given the example above, we want a relation $\mathfrak{R}_{a}|\phi^{\mathscr{B}}$ with $\mathfrak{R}_{a}|\phi^{\mathscr{B}}\mathscr{B}\mathscr{C}$, because $\phi$ holds at both of these bases, and $\mathfrak{R}_{a}|\phi^{\mathscr{B}}\mathscr{B}^{\prime}\mathscr{C}^{\prime}$, so that $(d)$ holds, and a relation $\mathfrak{R}_{a}|\phi^{\mathscr{B}^{\prime}}$ without $\mathfrak{R}_{a}|\phi^{\mathscr{B}^{\prime}}\mathscr{B}^{\prime}\mathscr{C}^{\prime}$, because they disagree on $\phi$, and $\mathfrak{R}_{a}|\phi^{\mathscr{B}^{\prime}}\mathscr{B}\mathscr{C}$, so that $(c)$, the converse of $(d)$, holds (see Figure 10).
$\mathscr{B}^{\prime}$$\phi$$\mathscr{C}^{\prime}$$\not{\phi}$$\mathscr{B}$$\phi$$\mathscr{C}$$\phi$$a$$a$ $\mathscr{B}^{\prime}$$\phi$$\mathscr{C}^{\prime}$$\not{\phi}$$\mathscr{B}$$\phi$$\mathscr{C}$$\phi$
Figure 10: Updating $\mathfrak{R}_{a}$ from Figure 9 by the announcement of $\phi$ results in different modal relations: $\mathfrak{R}_{a}|\phi^{\mathscr{B}}$ (left) and $\mathfrak{R}_{a}|\phi^{\mathscr{B^{\prime}}}$ (right).
As it turns out multiple modal relations can fill the bill and work as the $\mathfrak{R}_{A}|\phi$s we require and we do not have any reason to prefer one over the others. For that reason we treat announcement as a box-like operator that ranges over any such update. Note that it is not uncommon to treat announcements as box-operators, in Kripke semantics, however, there is exactly one modal update that corresponds to each announcement. This is a difference between the two semantic treatments of announcements.
[…]
Definition 14.
A base $\mathscr{C}$ is considered reachable from a base $\mathscr{B}$ by a set of relations $\mathfrak{R}_{A}^{\mathscr{B}}$ iff there are bases $\mathscr{C}_{1},\dots,\mathscr{C}_{n}$ and agents $a_{0},\dots a_{n}\in A$ s.t. $\mathfrak{R}_{a_{0}}\mathscr{B}\mathscr{C}_{1}$, $\mathfrak{R}_{a_{1}}\mathscr{C}_{1}\mathscr{C}_{2}$, $\dots$, $\mathfrak{R}_{a_{n}}\mathscr{C}_{n}\mathscr{C}$ (i.e., there is a way to get from $\mathscr{B}$ to $\mathscr{C}$ using some combination of the relations in $\mathfrak{R}_{A}^{\mathscr{B}}$).
We simultaneously define effective updates (Definition 15) and the support relation for PAL (Definition 16). However, we give the definitions separately to make them easier to reference. Note that, although the definition of effective updates makes use of the support relation (and vice-versa), the effective updates invoked in the definition of the support relation are on formulae of lesser height and so do not cause problems with the definitions.
As might be expected, the case for formulae of the form $[\phi]\psi$ depends on whether $\psi$ holds given effective updates by $\phi$. Accordingly, we define the support relation ,$\Vdash^{\phi}_{\mathscr{B},\mathfrak{R}_{A}}$ for after an announcement of $\phi$. We use $\Delta$ for a sequence of announcements. Maybe more surprisingly, the $K_{a}\phi$ case is the most changed. If we are evaluating $K_{a}\phi$, the condition remains unchanged from the modal condition in Definition 12. However, if we are evaluating $K_{a}\phi$ after an announcement, then we cannot just consider all modal relations that are effective updates of the initial relation. We also have to check all superset bases of our initial base. This is exactly because we are no longer considering a single modal relation, but rather multiple and so condition (d) of Definition 11 does not guarantee that all bases reachable by a superset of our initial base also have corresponding subset bases that are reachable from the initial base. Take the example as shown in Figure 10, even though for some $\mathfrak{R}_{a}$ we have $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$, we do not necessarily have $\mathfrak{R}^{\prime}_{a}\mathscr{B}^{\prime}\mathscr{C}^{\prime}$ for some corresponding $\mathfrak{R}^{\prime}_{a}$.
Note that as announcements only influence the relations we consider the case for propositional atoms, $\bot$, and implication remain unchanged from their $S5$ versions in Definition 12.
Definition 15 (Effective Updates).
Given a base $\mathscr{B}$, a set of relations $\mathfrak{R}_{A}$ and a formula $\phi$, $\mathfrak{R}^{\prime}_{A}$ is an effective update of $\mathfrak{R}_{A}$ by $\phi$ at $\mathscr{B}$ iff for all agents $a\in A$ and bases $\mathscr{C}$ and $\mathscr{D}$ reachable by $\mathfrak{R}_{A}$ from $\mathscr{B}$,
$\mathfrak{R}^{\prime}_{a}\mathscr{C}\mathscr{D}$ iff $\mathfrak{R}_{a}\mathscr{C}\mathscr{D}$, $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$, and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\phi$.
Updates that are not effective are called ineffective. We let $\mathfrak{R}_{A}|\Delta^{\mathscr{B}}$ denote the set of effective updates of $\mathfrak{R}_{A}$ by $\Delta$ at $\mathscr{B}$.
Definition 16 (Validity).
The support relation $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}$, where $\Delta$ is a (possibly empty) sequence of formulae, is defined as follows:
$\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}p&\mbox{iff}&\mbox{$p$ is in every set of basic sentences closed under $\mathscr{B}$}\\ &&\mbox{(i.e., iff $p\in\overline{\mathscr{B}}$)}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi\to\psi&\mbox{iff}&\mbox{$\phi\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\psi$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\bot&\mbox{iff}&\mbox{$\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}p$ for every basic sentence $p$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}K_{a}\phi&\mbox{iff}&\mbox{-- if$\Delta$ is empty, then for all $\mathscr{B}^{\prime}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{B}^{\prime}$, $\Vdash_{\mathscr{B}^{\prime},\mathfrak{R}_{A}}^{\Delta}\phi$}\\ &&\mbox{-- if $\Delta$ is non-empty, then for all $\mathscr{C}\supseteq\mathscr{B}$, $\mathfrak{R}_{A}^{\prime}\in\mathfrak{R}_{A}|\Delta^{\mathscr{C}}$,}\\ &&\mbox{\,\; and $\mathscr{C}^{\prime}$ s.t. $\mathfrak{R}^{\prime}_{a}\mathscr{C}\mathscr{C^{\prime}}$, $\Vdash_{\mathscr{C}^{\prime},\mathfrak{R}_{A}}^{\Delta}\phi$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}[\phi]\psi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\phi$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta,\phi}\psi$}\\ &&\\ \Gamma\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\psi$ for all $\psi\in\Gamma$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\phi$}\\ \end{array}$
We write $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ for $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$ with empty $\Delta$. A formula $\phi$ is valid, written as $\Vdash\phi$, iff $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ for all modal bases $\mathscr{B}$ and sets of $S5$-modal relations $\mathfrak{R}_{A}$.
To show that there always is at least one such effective update of $\mathfrak{R}_{A}$ by $\phi$ at a base $\mathscr{B}$ whenever $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$, we define a construction that results in such a relation.
Definition 17.
Given a set relations on bases $\mathfrak{R}_{A}$, let $|\mathfrak{R}_{A}|$ be the set of bases in the domains and ranges of the relations in $\mathfrak{R}_{A}$, that is
$\mathscr{B}\in|\mathfrak{R}_{A}|$ iff there exist $\mathfrak{R}_{a}\in\mathfrak{R}_{A}$ and $\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$ or $\mathfrak{R}_{a}\mathscr{C}\mathscr{B}$.
Additionally, let $\Delta(\mathfrak{R}_{A})$ be the set of all base rules of the bases in $|\mathfrak{R}_{A}|$, that is
$\Delta(\mathfrak{R}_{A})=\bigcup_{\mathscr{B}\in|\mathfrak{R}_{A}|}\mathscr{B}$
We define a construction of a relation after an announcement takes place. We do so by first fixing the relations between bases reachable by our initial base and then fixing the relations between subsets and supersets of those bases in a way that guarantees that the resulting relation is a modal relation.
Definition 18.
For all bases $\mathscr{B}$ and modal relations $\mathfrak{R}_{a}$, we define $\mathfrak{R}_{a}^{\mathscr{B}}=\mathfrak{R}_{a}.$
Given a set of agents $A$, we define a relation on the set of bases $S_{a}^{\mathscr{B}}$ to be the subset of $\mathfrak{R}_{a}$ restricted to bases reachable from $\mathscr{B}$ by $\mathfrak{R}_{A}$, that is to say,
$S_{a}^{\mathscr{B}}\mathscr{C}\mathscr{D}&\mbox{iff}&\mbox{$\mathscr{C}$ and $\mathscr{D}$ are reachable from $\mathscr{B}$ by $\mathfrak{R}_{A}^{\mathscr{B}}$ and $\mathfrak{R}_{a}^{\mathscr{B}}\mathscr{C}\mathscr{D}$.}\\ \end{array}$
Given this definition, we construct $\mathfrak{R}_{a}|\phi^{\mathscr{B}}$ in four steps:
Step 1: We update $S_{a}^{\mathscr{B}}$ by the public announcement of $\phi$.
$S_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{D}&\mbox{iff}&\mbox{$S_{a}^{\mathscr{B}}\mathscr{C}\mathscr{D}$ and $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ iff $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\phi$}\par\end{array}$
Additionally we take $S_{A}|\phi^{\mathscr{B}}$ to be the set of $S_{a}|\phi^{\mathscr{B}}$ for all agents $a\in A$.
Step 2: Here we eliminate those relations between bases that are no longer reachable from $\mathscr{B}$ by $S_{A}|\phi^{\mathscr{B}}$ as they are no longer relevant to $\mathscr{B}$ and can cause trouble with the modal relation conditions of Definition 11.
$S^{*}_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{D}&\mbox{iff}&\mbox{$S_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{D}$ and $\mathscr{C}$ and $\mathscr{D}$ are reachable from $\mathscr{B}$ by $S_{A}|\phi^{\mathscr{B}}$.}\par\end{array}$
Additionally we take $S^{*}_{A}|\phi^{\mathscr{B}}$ to be the set of $S^{*}_{a}|\phi^{\mathscr{B}}$ for all agents $a\in A$.
Step 3: We now want to add the proper subsets of the bases in $S^{*}_{A}|\phi^{\mathscr{B}}$ into the relation in a way that ensures that the modal relation conditions of Definition 11 are preserved. Informally, we keep the relations from $S^{*}_{A}|\phi^{\mathscr{B}}$ and, for the other bases, we establish a relation between $\mathscr{B}$ and $\mathscr{C}$ iff all supersets of $\mathscr{B}$ in $|S^{*}_{A}|\phi^{\mathscr{B}}|$ are connected to a superset of $\mathscr{C}$ in $S^{*}_{A}|\phi^{\mathscr{B}}$ and vice versa.
[…]
Again, we take $T_{A}|\phi^{\mathscr{B}}$ to be the set of $T_{a}|\phi^{\mathscr{B}}$ for all agents $a\in A$.
Step 4: Now we need to deal with the rest of the bases, that is the supersets of the bases in $|T_{A}|\phi^{\mathscr{B}}|$. Note that this includes all the inconsistent bases. Again we pick our relations between bases so that the modal relation conditions hold.
Like before we take the relations established in the previous step. We also add relations between all inconsistent bases. Analogously to step 2, we add relations between bases for which all subsets in $|T_{A}|\phi^{\mathscr{B}}|$ have appropriate relations in $T_{A}|\phi^{\mathscr{B}}$. The difference here is that we now also have to consider bases that contain rules outside of $\Delta(T_{A}|\phi^{\mathscr{B}})$. We handle those by only allowing relations between bases that have the same such new rules.
$R_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{D}&\mbox{iff}&\mbox{either $T_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{D}$}\\ &&\\ &&\mbox{or $\mathscr{C}$ and $\mathscr{D}$ are consistent and $\mathscr{C},\mathscr{D}\notin|T_{A}|\phi^{\mathscr{B}}|$ and, for all}\\ &&\mbox{$\mathscr{C}^{\prime}\subset\mathscr{C}$ s.t $\mathscr{C}^{\prime}\in|T_{A}|\phi^{\mathscr{B}}|$, there exists $\mathscr{D}^{\prime}\subset\mathscr{D}$ s.t. $T_{A}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$}\\ &&\mbox{and, for all $\mathscr{D}^{\prime}\subset\mathscr{D}$ s.t $\mathscr{D}^{\prime}\in|T_{A}|\phi^{\mathscr{B}}|$, there exists $\mathscr{C}^{\prime}\subset\mathscr{C}$ s.t. }\\ &&\mbox{$T_{A}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$ and$\mathscr{C}\setminus\Delta(T_{A}|\phi^{\mathscr{B}})=\mathscr{D}\setminus\Delta(T_{A}|\phi^{\mathscr{B}})$}\\ &&\\ &&\mbox{or $\mathscr{C}$ and $\mathscr{D}$ are inconsistent}\\ \end{array}$
Finally, we take $R_{A}|\phi^{\mathscr{B}}$ to be the set of $R_{a}|\phi^{\mathscr{B}}$ for all $a\in A$.
We now need to show that this is in fact results in $S5$ modal relations.
Lemma 6.
For any set modal relations $\mathfrak{R}_{A}$, $a\in A$, base $\mathscr{B}$ and formula $\phi$, $R_{a}|\phi^{\mathscr{B}}$ is a $S5$-modal relation.
Proof.
We have to show that the conditions of Definition 11 hold at $R_{a}|\phi^{\mathscr{B}}$.
To see that (a) and (b) still hold, simply note that all inconsistent bases are connected.
To show that condition (d) holds, take some consistent $\mathscr{C}$ and $\mathscr{D}$ s.t. $R_{A}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{D}$. We show that for any $\mathscr{C}^{\prime}\supseteq\mathscr{C}$ there is a $\mathscr{D}^{\prime}\supseteq\mathscr{D}$ with $R_{A}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$.
Before we begin we use a bit of notation: In the following we use $\mathscr{C}^{*}$ for any superset of $\mathscr{C}\in|S^{*}_{A}|\phi^{\mathscr{B}}|$. Additionally, let $\mathscr{C}^{-}=\mathscr{C}\cap\Delta(T_{A}|\phi^{\mathscr{B}})$ (i.e., the largest subset of $\mathscr{C}\in|T_{A}|\phi^{\mathscr{B}}|$).
[…]
Since $\mathfrak{R}_{a}$ is a modal relation with $\mathfrak{R}_{a}\mathscr{C}\mathscr{D}$ and $\mathfrak{R}_{a}\mathscr{E}\mathscr{F}$ and $\mathscr{C}^{\prime}$ is a subset of both $\mathscr{C}$ and $\mathscr{E}$, we know that there exists a $\mathscr{G}$ that is a subset of $\mathscr{D}$ and $\mathscr{F}$ s.t. $\mathfrak{R}_{a}\mathscr{C}^{\prime}\mathscr{G}$. Let $\mathscr{D}^{\prime}$ be the largest subset base $\mathscr{D}$ and $\mathscr{E}$ share.
Since, for all $\mathscr{C}^{*}$ and $\mathscr{D}^{*}$, $S^{*}_{a}|\phi^{\mathscr{B}}\mathscr{C}^{*}\mathscr{D}^{*}$ and $S^{*}_{a}|\phi^{\mathscr{B}}\mathscr{E}\mathscr{F}$, it follows, from step 3 of our construction, that $T_{a}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$. Which completes our proof by contradiction.
Now we consider the case in which $\mathscr{C}\notin|T_{A}|\phi^{\mathscr{B}}|$. If $\mathscr{C}^{\prime}\in|T_{A}|\phi^{\mathscr{B}}|$, this follows immediately by construction of $T_{a}|\phi^{\mathscr{B}}$ in step 2. Otherwise, we know that $\mathscr{C}^{-}\in|T_{A}|\phi^{\mathscr{B}}|$ and that $T_{a}|\phi^{\mathscr{B}}\mathscr{C}^{-}\mathscr{D}^{-}$, by step 4 of the construction. Similarly, we know that $\mathscr{C}^{{}^{\prime}-}\in|T_{A}|\phi^{\mathscr{B}}|$ and that there exists a $\mathscr{D}^{{}^{\prime}-}$ s.t. $T_{a}|\phi^{\mathscr{B}}(\mathscr{C}^{{}^{\prime}-}\mathscr{D}^{{}^{\prime}-}$. Let $\mathscr{D}^{\prime}=\mathscr{D}^{{}^{\prime}-}\cup(\mathscr{C^{\prime}}\setminus\Delta(T_{A}|\phi^{\mathscr{B}}))$ and, by steo 4, we can conclude $R_{a}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$. So, condition (d) holds.
As with condition (d), we consider two cases for condition (c). First, let $\mathscr{C}^{\prime}\in|T_{A}|\phi^{\mathscr{B}}|$. So, for every $\mathscr{C}^{*}$, there is a $\mathscr{D}^{*}$ s.t. $S_{a}^{*}|\phi^{\mathscr{B}}\mathscr{C}^{*}\mathscr{D}^{*}$. Specifically, this includes those $\mathscr{C}^{*}\supset\mathscr{C}^{\prime}$ and so there are those $\mathscr{D}^{\prime}$ that are subsets of the corresponding $\mathscr{D}^{*}$ and we have $T_{a}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$ and $\mathfrak{R}_{a}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$.
If $\mathscr{C}^{\prime}\notin|T_{A}|\phi^{\mathscr{B}}|$, we know there are $\mathscr{C}^{-}$ and $\mathscr{D}^{-}$ s.t. $T_{a}|\phi^{\mathscr{B}}\mathscr{C}^{-}\mathscr{D}^{-}$, by construction step 4, and, by the above, $\mathscr{C}^{{}^{\prime}-}$ and $\mathscr{D}^{{}^{\prime}-}$ s.t. $T_{a}|\phi^{\mathscr{B}}\mathscr{C}^{{}^{\prime}-}\mathscr{D}^{{}^{\prime}-}$. Let $\mathscr{D}^{\prime}=\mathscr{D}^{{}^{\prime}-}\cup(\mathscr{C^{\prime}}\setminus\Delta(T_{A}|\phi^{\mathscr{B}}))$. So, by our construction we have $R_{A}|\phi^{\mathscr{B}}\mathscr{C}^{\prime}\mathscr{D}^{\prime}$.
This suffices to show that $R_{a}|\phi^{\mathscr{B}}$ is a modal relation. What is left is to show that it is an equivalence relation.
Note that since $\mathfrak{R}_{a}$ is an equivalence relation, so is $S_{a}^{*}|\phi$.
To show that $T_{a}|\phi^{\mathscr{B}}$ is a reflexive it suffices to point that for any bases $\mathscr{C}$ and $\mathscr{C}^{*}$, we have $S_{a}^{*}|\phi^{\mathscr{B}}\mathscr{C}^{*}\mathscr{C}^{*}$ and so $T_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{C}$. To show that $R_{a}|\phi^{\mathscr{B}}$ is reflexive, it is only left to show that for any base $\mathscr{C}\notin T_{a}|\phi^{\mathscr{B}}$, we have $R_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{C}$. If $\mathscr{C}$ is inconsistent this is directly by construction. Otherwise note that we have $R_{a}|\phi^{\mathscr{B}}\mathscr{C}^{-}\mathscr{C}^{-}$ and so we also have $R_{a}|\phi^{\mathscr{B}}\mathscr{C}\mathscr{C}$.
The case for transitivity and Euclidean follow analogous reasoning. ∎
As we have shown the resulting relation is always a modal relation, however, we also need to show that the update is an effective update.
Lemma 7.
For any set of relation $\mathfrak{R}_{A}$ and formulae $\phi$ s.t. $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$, $R_{A}|\phi^{\mathscr{B}}$ is an effective update of $\mathfrak{R}_{A}$ at $\mathscr{B}$ by $\phi$, that is, $R_{A}|\phi^{\mathscr{B}}\in\mathfrak{R}_{A}|\phi^{\mathscr{B}}$.
Proof.
As shown in Lemma 6, $\mathfrak{R}_{A}|\phi^{\mathscr{B}}$ is a set of $S5$ modal relations. To show that the update was effective it suffices to point out that $S^{*}_{A}|\phi^{\mathscr{B}}$ fulfills the necessary criteria and the construction of $T_{A}|\phi^{\mathscr{B}}$ and $\mathfrak{R}_{A}|\phi^{\mathscr{B}}$ only alter proper sub- and supersets of the bases in $|S^{*}_{A}|\phi^{\mathscr{B}}|$ but not those bases itself. ∎
Corollary 1.
For any $\mathfrak{R}_{A}$, $\mathscr{B}$, and $\phi$ s.t. $\Vdash_{\mathscr{B},\mathscr{R}_{A}}\phi$, there is an effective update of $\mathfrak{R}_{A}$ by $\phi$ at $\mathscr{B}$.
We now move to show soundness and completeness for our system. We start by showing that the axioms and rules of $S5$ modal logic still hold as they did before adding announcements.
For this we need to adept some lemmas from the modal case. First we show that, even with our new updated relations, a formula that holds at a base will also hold at its supersets.
Lemma 8.
If $\Gamma\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$ and $\mathscr{B}\subseteq\mathscr{C}$, then $\Gamma\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\phi$.
Proof.
For empty $\Delta$, it suffices to point out that the only new case compared to Lemma 1 is the case that $\phi=[\psi]\chi$ which follows immediately from Definition 16.
The cases for non-empty $\Delta$ follow in exactly the same way for all cases except $\phi=K_{a}\psi$. Note that this case follows immediately from Definition 16 as well. ∎
We also show that every formula holds at inconsistent bases.
Lemma 9.
For all sets of $S5$-modal relations $\mathfrak{R}_{A}$, $\Delta$, and bases $\mathscr{B}$, if $\mathscr{B}$ is inconsistent, then $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$ for all $\phi$.
Proof.
We prove this by induction on the complexity of $\phi$. For $\phi=p$ and ϕ=italic-$\phi=\bot$ this follows immediately from the definition of inconsistent bases and the validity conditions in Definition 16, as they do not depend on the choice of the modal relation. For $\phi=\psi\to\chi$, we have $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\psi\to\chi$ iff $\phi\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\psi$. Since all $\mathscr{C}\supseteq\mathscr{B}$ s.t. $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\psi$ will also be inconsistent by Lemma 8, we know that $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\chi$ by the induction hypothesis.
For the next case let $\phi=K_{a}\psi$. By Lemma 8, all $\mathscr{C}\supseteq\mathscr{B}$ are also inconsistent and, for the case in which $\Delta$ is non-empty, so are all $\mathscr{D}$ for which there is a $\mathfrak{R}^{\prime}\in\mathfrak{R}_{A}|\Delta^{\mathscr{C}}$ with $\mathfrak{R}_{a}\mathscr{C}\mathscr{D}$ and we have $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\Delta}\psi$ by the induction hypothesis. For the case in which $\Delta$ is empty, it suffices to point out that, due to condition $(a)$, all bases $\mathscr{X}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{X}$ are also inconsistent and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\Delta}\psi$ by the induction hypothesis.
Finally, let $\phi=[\psi]\chi$. For any $\mathscr{C}\supseteq\mathscr{B}$, if we have $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\bot$, then we also have $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta,\phi}\bot$ by the validity conditions of $\bot$. So, by induction hypothesis, $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta,\phi}\chi$ and $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}[\psi]\chi$. ∎
Given this, we can now look at maximally-consistent bases which will turn out to be helpful for proving the completeness of our semantics.
We show that whether a formula holds at a maximally-consistent base can be established without referencing any of its supersets.
Lemma 10.
For any set of $S5$-modal relations $\mathfrak{R}_{A}$, $\Delta$, and maximally-consistent base $\mathscr{B}$, the following hold:
$\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\bot&&\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi\to\psi&\mbox{iff}&\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi\mbox{ or }\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\psi\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}K_{a}\phi&\mbox{iff}&\mbox{for all $\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{C},\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\phi$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}[\phi]\psi&\mbox{iff}&\mbox{ $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$ implies $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta,\phi}\psi$}\\ \end{array}$
Proof.
The $\bot$-cases follow from directly from consistency.
For the other cases it suffices to point out that any $\mathscr{B}^{\prime}\supset\mathscr{B}$ is inconsistent and so by Lemma 9 any formula is supported at them given any $\mathfrak{R}_{A}$ or $\mathfrak{R}_{A}|\Delta^{\mathscr{B}^{\prime}}$. ∎
From this, it follows that the law of excluded middle holds on maximally-consistent bases: for every formula $\phi$, either $\phi$ or ϕitalic-$\phi\to\bot$.
Lemma 11.
For every formula $\phi$, maximally-consistent base $\mathscr{B}$, $S5$-modal relations $\mathfrak{R}_{A}$, and $\Delta$, either $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$ or $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi\to\bot$.
Proof.
We assume $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$ and show $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi\to\bot$.
By Lemma 10, $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi\to\bot$ iff if $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$, then $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\bot$. So, since we assumed $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$, the antecedent is false and so $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi\to\bot$. ∎
There is one more important property of maximally-consistent bases that we require for the soundness proof. If a formula does not hold at a base, then that base can be extended to a maximally-consistent base on which the formula does not hold.
To show this we need two more results about the compositionality of announcements.
Lemma 12.
For any formulae $\phi$ and $\psi$, base $\mathscr{B}$, and set of $S5$-modal relations $\mathfrak{R}_{A}$, $\mathfrak{R}_{A}|(\phi\wedge[\phi]\psi)^{\mathscr{B}}=\mathfrak{R}_{A}|(\phi;\psi)^{\mathscr{B}}$.
Proof.
We show this by showing that for all $\mathscr{C},\mathscr{D}$ reachable from $\mathscr{B}$ by $\mathfrak{R}_{a}$ these two updates give us the same result. i.e. $S^{*}_{a}|(\phi\wedge[\phi]\psi)^{\mathscr{B}}=S^{*}_{a}|(\phi;\psi)^{\mathscr{B}}$ (as defined in Definition 18). Note that, if $S_{a}^{\mathscr{B}}\mathscr{C}\mathscr{D}$, then $S_{a}^{\mathscr{B}}=S_{a}^{\mathscr{C}}=S_{a}^{\mathscr{D}}$.
For all $\mathscr{C},\mathscr{D}$, $S_{a}|(\phi\wedge[\phi]\psi)^{\mathscr{B}}\mathscr{C}\mathscr{D}$ iff $S_{a}^{\mathscr{B}}\mathscr{C}\mathscr{D}$ and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}^{\mathscr{D}}}\phi\wedge[\phi]\psi$ iff $S_{a}^{\mathscr{B}}\mathscr{C}\mathscr{D}$, $\Vdash_{\mathscr{D},\mathfrak{R}_{A}^{\mathscr{D}}}\phi$ and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}|\phi^{\mathscr{D}}}\psi$. Similarly, $S_{a}|{\phi;\psi}^{\mathscr{B}}\mathscr{C}\mathscr{D}$ iff $S_{a}^{\mathscr{B}}\mathscr{C}\mathscr{D}$, $\Vdash_{\mathscr{D},\mathfrak{R}_{A}^{\mathscr{D}}}\phi$ and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}|\phi^{\mathscr{D}}}\psi$.
Since $S_{a}|(\phi\wedge[\phi]\psi)^{\mathscr{B}}=S_{a}|(\phi;\psi)^{\mathscr{B}}$, we also know that $S^{*}_{a}|(\phi\wedge[\phi]\psi)^{\mathscr{B}}=S^{*}_{a}|(\phi;\psi)^{\mathscr{B}}$. ∎
Corollary 2.
For every string of announcements $\Delta=\phi_{1};\dots;\phi_{n}$, there is a single formula $\phi(\Delta)$ s.t., for every $\mathscr{B}$, $\mathfrak{R}_{a}|\phi(\Delta)^{\mathscr{B}}=\mathfrak{R}_{a}|\Delta^{\mathscr{B}}$. Specifically, for non-empty $\Delta$, $\phi(\Delta)=\phi_{1}\wedge[\phi_{1}](\phi_{2}\wedge[\phi_{2}](\phi_{3}\wedge\dots\wedge[\phi_{n-1}]\phi_{n}))$. For empty $\Delta$, $\phi(\Delta)$ can be any tautology.
Now we can show the lemma we initially wanted.
Lemma 13.
For any string of announcements $\Delta$, formula $\phi$, base $\mathscr{B}$ and a set of $S5$-model relations $\mathfrak{R}_{A}$, if $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi(\Delta)$ and $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\phi$, then there exists a maximally-consistent base $\mathscr{B}^{*}\supseteq\mathscr{B}$ with $\nVdash_{\mathscr{B}^{*},\mathfrak{R}_{A}}^{\Delta}\phi$.
Proof.
We proof these two results simultaneously by induction on the complexity of $\phi$. The cases for $\phi=p$ and ϕ=italic-$\phi=\bot$ are straightforward, as the choice of $\mathfrak{R}$ is irrelevant to them.
Note that for for empty $\Delta$ the cases for $\phi=K_{a}\psi$ and $\phi=\psi\to\chi$ follow immediately from Lemma 5. We proceed to show these cases for non-empty $\Delta$.
For $\phi=\psi\to\chi$, we start with $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}\psi\to\chi$. So, there is a $\mathscr{C}\supseteq\mathscr{B}$ s.t. $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\psi$ and $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\chi$. By induction hypothesis, there is a maximally-consistent $\mathscr{C}^{*}\supseteq\mathscr{C}$ with $\nVdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}^{\Delta}\chi$ and, since $\mathscr{C}^{*}\supseteq\mathscr{C}$, $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}^{\Delta}\psi$. So, $\nVdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}^{\Delta}\psi\to\chi$.
The interesting case is the one in which $\phi=K_{a}\psi$. Given $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}K_{a}\psi$, there are $\mathscr{X}\supseteq\mathscr{B}$, $\mathscr{C}$ and $\mathfrak{R}^{\prime}_{A}\in\mathfrak{R}_{A}|\Delta^{\mathscr{X}}$ s.t. $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\psi$, by induction hypothesis, there is a maximally-consistent $\mathscr{C}^{*}\supseteq\mathscr{C}$ with $\nVdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\psi$. From $\mathfrak{R}^{\prime}_{a}\mathscr{X}\mathscr{C}$, we know that $\mathfrak{R}_{a}\mathscr{X}\mathscr{C}$ and, by symmetry, $\mathfrak{R}_{a}\mathscr{C}\mathscr{X}$. Since $\mathfrak{R}_{a}$ is a modal relation, we know that, by condition (c) of Definition 11, there is a $\mathscr{X}^{\prime}\supseteq\mathscr{X}$ s.t. $\mathfrak{R}_{a}\mathscr{C}^{*}\mathscr{X}^{\prime}$ and, again by symmetry, that $\mathfrak{R}_{a}\mathscr{X}^{\prime}\mathscr{C}^{*}$. Again by (c), for every maximally-consistent $\mathscr{X}^{*}\supseteq\mathscr{X}^{\prime}$, $\mathfrak{R}_{a}\mathscr{X}^{*}\mathscr{C}^{*}$, because $\mathscr{C}^{*}$ does not have any consistent superset bases other than itself. Now we only need to show that the relation between those two bases still holds for some relation $\mathfrak{R}^{\prime\prime}_{A}\in\mathfrak{R}_{A}|\Delta^{\mathscr{X}^{*}}$.
Since, by assumption, $\Vdash_{\mathscr{X},\mathfrak{R}_{A}}\phi(\Delta)$ and $\mathscr{X}^{*}\supseteq\mathscr{X}$, we also have $\Vdash_{\mathscr{X}^{*},\mathfrak{R}_{A}}\phi(\Delta)$. Similarly, because of $\mathfrak{R}^{\prime}_{a}\mathscr{X}\mathscr{C}$, we know that $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi(\Delta)$ and so $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\phi(\Delta)$. So, we have $\mathfrak{R}^{\prime\prime}_{a}\mathscr{X}^{*}\mathscr{C}^{*}$ and, finally, $\nVdash_{\mathscr{X}^{*},\mathfrak{R}_{A}}^{\Delta}K_{a}\psi$.
Finally, consider the case that $\phi=[\psi]\chi$. Take $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}^{\Delta}[\psi]\chi$, so there is a $\mathscr{C}\supseteq\mathscr{B}$ s.t. $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta}\psi$ but $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}^{\Delta,\phi}\chi$. So, by induction hypothesis, there is a maximally-consistent such $\mathscr{C}^{*}\supseteq\mathscr{C}$ and we are done. ∎
We can now finally show that the rules and axioms of modal logic hold for our base-extension semantics of PAL. We break this up to make it easier on the reader and start with the classical axioms and modus ponens.
Lemma 14.
The rules and axioms of $S5$ modal logic are valid on the base-extension semantics for PAL.
  • (1) $\phi\Vdash\psi\to\phi$
  • (2) $\phi\to(\psi\to\chi)\Vdash(\phi\to\psi)\to(\phi\to\chi)$
  • (3) $(\phi\to\bot)\to(\psi\to\bot)\Vdash\psi\to\phi$
  • (MP) If $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to\psi$ and $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$, then $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi$
Proof.
We start by showing (MP) as it is helpful in establishing the other results. Simply note that because $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to\psi$, we have for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\psi$. By assumption, $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$ and so $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi$.
By Definition 16, for (1) we need to show that for any $\mathscr{B}$, if $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$, then $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi\to\phi$. By Lemma 8, for any $\mathscr{C}\supseteq\mathscr{B}$ s.t. $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\psi$ we also have $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ and so $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi\to\phi$.
For (2) we follow the same strategy of showing that if the formula on the left of the turnstyle holds at a base, so does the right. To show that $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}(\phi\to\psi)\to(\phi\to\chi)$, we need to show that for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi\to\psi$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi\to\chi$. Take some $\mathscr{D}\supseteq\mathscr{C}$ s.t. $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\phi\to(\psi\to\chi)$, $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\phi\to\psi$ and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\phi$. By multiple application of (MP), we get $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\chi$ and so, $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi\to\chi$ and, finally, $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}(\phi\to\psi)\to(\phi\to\chi)$.
For (3) we take some $\mathscr{B}$ with $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}(\phi\to\bot)\to(\psi\to\bot)$. To show that $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi\to\phi$, we show that for all $\mathscr{C}\supseteq\mathscr{B}$ s.t. $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\psi$ we also have $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$. Assume $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ for contradiction. By Lemma 11, all maximally-consistent $\mathscr{C}^{*}\supseteq\mathscr{C}$ either $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\phi$ or $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\phi\to\bot$. By Lemma 8, we know that $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}(\phi\to\bot)\to(\psi\to\bot)$ and $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\psi$. So, if $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\phi\to\bot$, then, by (MP), $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\psi\to\bot$ and $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}\bot$. This is a contradiction as $\mathscr{C}^{*}$ is, by assumption, consistent. So, $\phi$ holds at all maximally-consistent superset bases of $\mathscr{C}$ and so, by the contrapositive of Lemma 13, $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ and so $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\psi\to\phi$. ∎
This concludes the classical portion. We continue with the modal axioms and (modal) (NEC).
Lemma 15.
The rules and axioms of $S5$ modal logic are valid on the base-extension semantics for PAL.
  • (K) $K_{a}(\phi\to\psi)\Vdash K_{a}\phi\to K_{a}\psi$
  • (T) $K_{a}\phi\Vdash\phi$
  • (4) $K_{a}\phi\Vdash K_{a}K_{a}\phi$
  • (5) $\neg K_{a}\phi\Vdash K_{a}\neg K_{a}\phi$
  • (NEC) If $\Vdash\phi$, then $\Vdash K_{a}\phi$
Proof.
For (K), we show the contrapositive and assume $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\phi\to K_{a}\psi$ for an arbitrary $\mathscr{B}$ and $\mathfrak{R}_{A}$. So, there is a $\mathscr{C}\supseteq\mathscr{B}$ s.t. $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}K_{a}\phi$ and $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}K_{a}\psi$. So, there are $\mathscr{D}\supseteq\mathscr{C}$ and $\mathscr{E}$ s.t. $\mathfrak{R}_{a}\mathscr{D}\mathscr{E}$ with $\Vdash_{\mathscr{E},\mathfrak{R}_{A}}\phi$ but $\nVdash_{\mathscr{E},\mathfrak{R}_{A}}\psi$ and so $\nVdash_{\mathscr{E},\mathfrak{R}_{A}}\phi\to\psi$ and $\nVdash_{\mathscr{D},\mathfrak{R}_{A}}K_{a}(\phi\to\psi)$. Finally, as $\mathscr{D}\supseteq\mathscr{B}$, $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}(\phi\to\psi)$.
For (T), take some $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\phi$. By the reflexivity of $\mathfrak{R}_{a}$, $\mathfrak{R}_{A}\mathscr{B}\mathscr{B}$ and so $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi$.
For (4), we, without loss of generality, assume some $\mathscr{B}$ and $\mathfrak{R}_{A}$ s.t. $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\phi$. For contradiction, we also assume $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}K_{a}\phi$. So, there are $\mathscr{C}\supseteq\mathscr{B}$ and $\mathscr{D}$ s.t. $\mathfrak{R}_{a}\mathscr{C}\mathscr{D}$ and $\nVdash_{\mathscr{D},\mathfrak{R}_{A}}K_{a}\phi$. This means there have to be $\mathscr{E}\supseteq\mathscr{D}$ and $\mathscr{F}$ s.t. $\mathfrak{R}_{a}\mathscr{E}\mathscr{F}$ and $\nVdash_{\mathscr{F},\mathfrak{R}_{A}}\phi$. By Definition 11 $(d)$, we know there is a $\mathscr{G}\subseteq\mathscr{F}$ and $\mathfrak{R}_{a}\mathscr{D}\mathscr{G}$ and, by transitivity, $\mathfrak{R}_{a}\mathscr{C}\mathscr{G}$. Obviously, $\nVdash_{\mathscr{G},\mathfrak{R}_{A}}\phi$ and so $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\phi$, which contradicts our assumption.
For (5), we assume, without loss of generality, some $\mathscr{B}$ and $\mathfrak{R}_{A}$ s.t. $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\neg K_{a}\phi$ and, for contradiction, $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\neg K_{a}\phi$. So, there is $\mathscr{C}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{C}$ and $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}\neg K_{a}\phi$. So, there is a $\mathscr{D}\supseteq\mathscr{C}$ s.t $\mathscr{D}$ is consistent and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}K_{a}\phi$.
From $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\neg K_{a}\phi$ it follows that for all consistent $\mathscr{E}\supseteq\mathscr{B}$, we have $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\phi$. So there is a $\mathscr{F}$ s.t. $\mathfrak{R}_{a}\mathscr{E}\mathscr{F}$ and $\nVdash_{\mathscr{F},\mathfrak{R}_{A}}\phi$.
By symmetry, we have $\mathfrak{R}_{a}\mathscr{C}\mathscr{B}$ and so, by condition (c) of Definition 11, for one of these $\mathscr{E}\supseteq\mathscr{B}$ we have $\mathfrak{R}_{a}\mathscr{D}\mathscr{E}$ and, again by symmetry, $\mathfrak{R}_{a}\mathscr{E}\mathscr{D}$. By Euclidean and $\mathfrak{R}_{a}\mathscr{E}\mathscr{F}$, we have $\mathfrak{R}_{a}\mathscr{D}\mathscr{F}$ which gives us our contradiction: We have $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}K_{a}\phi$ and $\nVdash_{\mathscr{F},\mathfrak{R}_{A}}\phi$.
We show (NEC) via the contrapositive. If $\nVdash K_{a}\phi$, then $\nVdash\phi$. Suppose $\nVdash K_{a}\phi$. So, there is a $\mathscr{B}$ and a $\mathfrak{R}_{A}$ s.t. $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}K_{a}\phi$. Therefore, there are $\mathscr{C}\supseteq\mathscr{B}$ and $\mathscr{D}$ s.t. $\mathfrak{R}_{a}\mathscr{C}\mathscr{D}$ and $\nVdash_{\mathscr{D},\mathfrak{R}_{A}}\phi$ and, so, $\nVdash\phi$. ∎
Before showing the axioms of PAL, we show Lemma 16. It is used in the proof of announcement and knowledge.
Lemma 16.
For all $\mathscr{B}$, $\mathfrak{R}_{A}$, and $\phi$ the following are equivalent:
  1. 1.
    For all $\mathscr{C}\supseteq\mathscr{B}$, $\mathfrak{R}^{\prime}_{A}\in\mathfrak{R}_{A}|\phi^{\mathscr{C}}$ and $\mathscr{D}$ s.t. $\mathfrak{R}^{\prime}_{a}\mathscr{C}\mathscr{D}$, $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi}\psi$.
  2. 2.
    For all $\mathscr{X}$ s.t. $\mathfrak{R}_{a}\mathscr{B}\mathscr{X}$ and $\mathscr{Y}\supseteq\mathscr{X}$, if $\Vdash_{\mathscr{Y},\mathfrak{R}_{A}}\phi$, then $\Vdash_{\mathscr{Y},\mathfrak{R}_{A}}^{\phi}\psi$.
Proof.
We show that any $\mathscr{D}$ in the former corresponds to some $\mathscr{Y}$ in the latter and vice versa.
We know that $\mathfrak{R}^{\prime}\mathscr{C}\mathscr{D}$, iff there is a $\mathscr{X}\subseteq\mathscr{D}$ s.t. $\mathfrak{R}\mathscr{B}\mathscr{X}$ and $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\phi$ since $\mathfrak{R}_{A}$ is a set modal relations and $\mathfrak{R}_{A}^{\prime}$ is an update of it by $\mathscr{\phi}$ at $\mathscr{C}$. For the bottom-to-top direction, it suffices to point out that from this it follows that the set of all $\mathscr{Y}$ (as in the third line of the proof) includes all $\mathscr{D}$. For the top-to-bottom direction, note that, for any $\mathscr{Y}$, there is a corresponding $\mathscr{C}$ s.t. $\mathfrak{R}_{A}\mathscr{C}\mathscr{Y}$, as $\mathfrak{R}_{A}$ is a set of modal relations, and since $\Vdash_{\mathscr{Y},\mathfrak{R}_{A}}\phi$ there is a corresponding $\mathfrak{R}_{A}^{\prime\prime}$ that is an effective update of $\mathfrak{R}_{A}$ by $\phi$ at $\mathscr{Y}$ s.t. $\mathfrak{R}^{\prime\prime}_{A}\mathscr{C}\mathscr{Y}$ and so every $\mathscr{Y}$ is also a $\mathscr{D}$. See Figure 11 for an illustration of this. ∎
$\mathscr{C}$$\mathscr{D}$$\mathscr{E}$$\phi$$\mathfrak{R}^{\prime}$Leftrightarrow $\mathscr{C}$$\mathscr{X}$$\mathscr{Y}$If $\phi$, then $\psi$$\mathfrak{R}$
Figure 11: Illustration of steps two (left) and three (right) of the proof of Announcement and Knowledge. Note that $\mathfrak{R}^{\prime}\mathscr{D}\mathscr{E}$ only if $\Vdash_{\mathscr{E},\mathfrak{R}^{\prime}}\phi$.
Given that we can now show that the new axioms and rules specific to PAL also hold for our system.
Lemma 17.
The axioms and rules of PAL are valid on the base-extension semantics for PAL.
$[\phi]p\leftrightarrow\phi\to p&\mbox{atomic permanence}\\ [\phi]\bot\leftrightarrow\phi\to\bot&\mbox{announcement and bot}\\ [\phi](\psi\to\chi)\leftrightarrow[\phi]\psi\to[\phi]\chi&\mbox{announcement and implication}\\ [\phi]K_{a}\psi\leftrightarrow\phi\to K_{a}[\phi]\psi&\mbox{announcement and knowledge}\\ [\phi][\psi]\chi\leftrightarrow[\phi\wedge[\phi]\psi]\chi&\mbox{announcement composition}\\ \mbox{If $\psi$, then $[\phi]\psi$}&\mbox{necessitation of announcements}\end{array}$
Proof.
$[\phi]p\leftrightarrow\phi\to p$
By definition of validity, $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}[\phi]p$ iff for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$, then $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}^{\phi}p$. Similarly, $\Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to p$ iff for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}p$. It suffices to point out that $\Delta$ has no bearing on the support of propositional $p$:
$[\phi]\bot\leftrightarrow\phi\to\bot$
The case for $\bot$ works analogously to the case for $p$. Simply note that $\bot$ holds iff all $p$ hold. So, this follows immediately from the result for atomic permanence:
$[\phi](\psi\to\chi)\leftrightarrow[\phi]\psi\to[\phi]\chi$
This case is done by simple unpacking of the definitions and moving the precondition $\phi$:
$\Vdash_{\mathscr{B},\mathfrak{R}_{A}}[\phi](\psi\to\chi)&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\phi}\psi\to\chi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then for all $\mathscr{D}\supseteq\mathscr{C}$, if $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi}\psi$,}\\ &&\mbox{then $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi}\chi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ implies that for all $\mathscr{D}\supseteq\mathscr{C}$,}\\ &&\mbox{$\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi}\psi$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ implies that for all $\mathscr{E}\supseteq\mathscr{C}$,}\\ &&\mbox{$\Vdash_{\mathscr{E},\mathfrak{R}_{A}}^{\phi}\chi$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}[\phi]\psi\to[\phi]\chi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}[\phi]\psi$,}\\ &&\mbox{then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}[\phi]\chi$}\par\end{array}$
For the steps between the second and third line, we simply moved $\Vdash_{\mathscr{C},\mathfrak{R}_{A}^{\mathscr{C}}}\phi$ and the universal quantification over the supersets of $\mathscr{C}$ between inside or outside the implication.
$[\phi]K_{a}\psi\leftrightarrow\phi\to K_{a}[\phi]\psi$
This case is also proven straightforwardly by unpacking and moving around the precondition $\phi$. The only exception is the step between the second and third line which holds by Lemma 16.
$\Vdash_{\mathscr{B},\mathfrak{R}_{A}}[\phi]K_{a}\psi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\phi}K_{a}\psi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then for all $\mathscr{D\supseteq\mathscr{C}}$, $\mathscr{E}$ and $\mathfrak{R}_{A}^{\prime}\in\mathfrak{R}_{A}|\phi^{\mathscr{D}}$,}\\ &&\mbox{s.t. $\mathfrak{R}^{\prime}_{a}\mathscr{D}\mathscr{E}$, $\Vdash_{\mathscr{E},\mathfrak{R}_{A}}^{\phi}\psi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then for all $\mathscr{X}$ s.t. $\mathfrak{R}_{a}\mathscr{C}\mathscr{X}$ and $\mathscr{Y}\supseteq\mathscr{X}$, if $\Vdash_{\mathscr{Y},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then $\Vdash_{\mathscr{Y},\mathfrak{R}_{A}}^{\phi}\psi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then for all $\mathscr{X}$ s.t. $\mathfrak{R}_{a}\mathscr{C}\mathscr{X}$, $\Vdash_{\mathscr{X},\mathfrak{R}_{A}}[\phi]\psi$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}\phi\to K_{a}[\phi]\psi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}K_{a}[\phi]\psi$}\\ \end{array}$
$[\phi][\psi]\chi\leftrightarrow[\phi\wedge[\phi]\psi]\chi$
This case, again, is proven by unpacking and because $\mathfrak{R}_{A}|\phi\wedge[\phi]\psi^{\mathscr{D}}=\mathfrak{R}_{A}|(\phi;\psi)^{\mathscr{D}}$ by Lemma 12:
$\Vdash_{\mathscr{B},\mathfrak{R}_{A}}[\phi][\psi]\chi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$,}\\ &&\mbox{then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\phi}[\psi]\chi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\phi}\phi$, then for all $\mathscr{D}\supseteq\mathscr{C}$,}\\ &&\mbox{ if $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi}\psi$, then $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi,\psi}\chi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$, then for all $\mathscr{D}\supseteq\mathscr{C}$,}\\ &&\mbox{if $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi}\psi$, then $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi\wedge[\phi]\psi}\chi$}\\ &\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ and for all $\mathscr{D}\supseteq\mathscr{C}$,}\\ &&\mbox{$\Vdash_{\mathscr{D},\mathfrak{R}_{A}}\phi$ implies $\Vdash_{\mathscr{D},\mathfrak{R}_{A}}^{\phi}\psi$, then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\phi\wedge[\phi]\psi}\chi$}\\ \Vdash_{\mathscr{B},\mathfrak{R}_{A}}[\phi\wedge[\phi]\psi]\chi&\mbox{iff}&\mbox{for all $\mathscr{C}\supseteq\mathscr{B}$, if $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi\wedge[\phi]\psi$,}\\ &&\mbox{then $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}^{\phi\wedge[\phi]\psi}\chi$}\\ \end{array}$
If $\psi$, then $[\phi]\psi$
We show this by contrapositive: In order for $\nVdash_{\mathscr{B},\mathfrak{R}_{A}}[\phi]\psi$, there has to be a $\mathscr{C}\supseteq\mathscr{B}$ with $\Vdash_{\mathscr{C},\mathfrak{R}_{A}}\phi$ and $\nVdash_{\mathscr{C},\mathfrak{R}_{A}}^{\phi}\psi$. By Lemma 13, there is a maximally-consistent $\mathscr{C}^{*}\supseteq\mathscr{C}$ s.t. $\nVdash_{\mathscr{C}^{*},\mathfrak{R}_{A}|\phi^{\mathscr{C}^{*}}}\psi$. Take some relation $\mathfrak{R}^{\prime}_{A}\in\mathfrak{R}_{A}^{\mathscr{C}^{*}}|\phi$. By Lemma 6, we know that $\mathfrak{R}^{\prime}_{a}$ is a $S5$-modal relation. From Lemma 10, it follows straightforwardly that for all formulae $\chi$, $\Vdash_{\mathscr{C}^{*},\mathfrak{R}_{A}}^{\phi}\chi$ iff $\Vdash_{\mathscr{C}^{*},\mathfrak{R}^{\prime}_{A}}\chi$. So we have $\nVdash_{\mathscr{C}^{*},\mathfrak{R}_{A}^{\prime}}\psi$, which contradicts our assumption that $\psi$ is valid. ∎
We can now proof the soundness of our system.
Theorem 2.
For all $\phi$, $\vdash\phi$ implies $\Vdash\phi$.
Proof.
This follows directly from Lemmas 14, 15 and 17. ∎
We establish completeness in a known way for PAL by translation between PAL and S5. We show that for every formula in PAL there is an equivalent formula in S5. The result then follows from the Completeness of S5 and the Soundness of PAL just shown. The steps of this proof are taken from [4]. We start by defining the necessary translation function based on the axioms for PAL.
Definition 19.
The translation $t:PAL\to S5$ is defined as follows:
$t(p)&=&p\\ t(\bot)&=&\bot\\ t(\phi\to\psi)&=&t(\phi)\to t(\psi)\\ t(K_{a}\phi)&=&K_{a}t(\phi)\\ t([\phi]p)&=&t(\phi\to p)\\ t([\phi]\bot)&=&t(\phi\to\bot)\\ t([\phi](\psi\to\chi))&=&t([\phi]\psi\to[\phi]\chi)\\ t([\phi]K_{a}\psi)&=&t(\phi\to K_{a}[\phi]\psi)\\ t([\phi][\psi]\chi)&=&t([\phi\wedge[\phi]\psi]\chi)\\ \end{array}$
This translation is used for an inductive proof to show that for every formula in PAL there is an equivalent formula in $S5$. To do this we need a complexity measure to show that such an induction concludes.
Definition 20.
The complexity $c:PAL\to\mathbb{N}$ is defined as follows:
$c(p)&=&1\\ c(\bot)&=&1\\ c(\phi\to\psi)&=&1+max(c(\phi),c(\psi))\\ c(K_{a}\phi)&=&1+c(\phi)\\ c([\phi]\psi)&=&(2+c(\phi))\cdot c(\psi)\\ \end{array}$
Before we go to the necessary complexity results, here are some helpful preliminary results based on the definitions of $\neg$ and $\wedge$ from our primary connectives.
Lemma 18.
The following equations hold:
  1. 1.
    $c(\neg\phi)=1+c(\phi)$
  2. 2.
    $c(\phi\wedge\psi)=3+max(1+c(\phi);c(\psi)$
Proof.
For the first equation,$\neg\phi$ is shorthand for ϕitalic-$\phi\to\bot$. We have $c(\phi\to\bot)=1+max(c(\phi);c(\bot))$ and $c(\bot)=1$. Depending on $\phi$, $c(\phi)$ is at least $1$ and so $max$ never has to pick $c(\bot)$ and so $c(\neg\phi)=1+c(\phi)$.
[…]
Given this measure the following hold:
Lemma 19.
The following hold:
  1. 1.
    $c(\psi)\geq c(\phi)$ if $\phi\in Sub(\psi)$
  2. 2.
    $c([\phi]p)>c(\phi\to p)$
  3. 3.
    $c([\phi]\bot)>c(\phi\to\bot)$
  4. 4.
    $c([\phi](\psi\to\chi))>c([\phi]\psi\to[\phi]\chi)$
  5. 5.
    $c([\phi]K_{a}\psi)>c(\phi\to K_{a}[\phi]\psi)$
  6. 6.
    $c([\phi][\psi]\chi)>c([\phi\wedge[\phi]\psi]\chi)$.
Proof.
The first result follows immediately from the definition of our complexity measure in Definition 20.
The second and third result follow in the same way as $c(p)=c(\bot)=1$. So, we only show it for $\bot$. $c(\phi\to\bot)=1+c(\phi)$ as in the proof of Lemma 18. $c([\phi]\bot)=(2+c(\phi))\times 1$. Obviously, $2+c(\phi)>1+c(\phi)$.
For the fourth result, $c([\phi]\psi\to\chi)=(2+c(\phi))\times(1+max(c(\psi);c(\chi)))$ and $c([\phi]\psi\to[\phi]\chi)=1+max(c([\phi]\psi;[\phi]\chi))=1+max(((2+c(\phi))\times c(\psi));((2+c(\phi))\times c(\chi)))=1+((2+c(\phi))\times max(c(\psi);c(\chi)))$. The result $(2+c(\phi))\times(1+max(c(\psi);c(\chi)))\geq 1+((2+c(\phi))\times max(c(\psi);c(\chi)))$ follows because $(2+c(\phi)$ is greater than $1$.
For the fifth result, $c([\phi]K_{a}\psi)=(2+c(\phi))\times(1+c(\psi))$ and $c(\phi\to K_{a}[\phi]\psi)=1+max(c(\phi);1+((2+c(\phi))\times c(\psi)))$. Obviously $c(\phi)$ cannot be greater than $1+((2+c(\phi))\times c(\psi))$ and so, we simplify to $2+((2+c(\phi))\times c(\psi))$. To show the result it suffices to point out that $2$ is less than $2+c(\phi)$.
For the final result of this lemma, $c([\phi][\psi]\chi)=(2+c(\phi))\times((2+c(\psi))\times c(\chi)$ and $c([\phi\wedge[\phi]\psi]\chi)=(2+c(\phi\wedge[\phi]\psi))\times c(\chi)=(2+3+max(1+c(\phi);c([\phi]\psi)))\times c(\chi)$. Note that $1+c(\phi)$ is never greater than $c([\phi]\psi)$ and so, we can simplify again to $(5+c([\phi]\psi))\times c(\chi)=(5+((2+c(\phi))\times c(\psi))\times c(\chi)$. If we compare $(2+c(\phi))\times((2+c(\psi))\times c(\chi)$ and $(5+((2+c(\phi))\times c(\psi))\times c(\chi)$, the former is greater if and only if $(2\times(2+c(\phi)\times c(\chi))$ is greater than $5$. This obviously holds. ∎
This allows us to establish the following result:
Lemma 20.
For all formulas $\phi$, $\Vdash\phi\leftrightarrow t(\phi)$.
Proof.
As mentioned above, we establish this by induction on the complexity of $\phi$.
For the base case note that $t(p)=p$. It is trivial to show that $\Vdash p\leftrightarrow p$.
All other case follow from the induction hypothesis and the corresponding case of Definition 20 and Lemma 19. ∎
From this and the completeness of $S5$ the following theorem follows immediately.
Theorem 3.
For all formulas of $PAL$, $\Vdash\phi$ implies $\vdash\phi$.
Proof.
Assume $\Vdash\phi$. By Theorem 2 and Lemma 20, we get $\vdash\phi\leftrightarrow t(\phi)$. Since $t(\phi)$ does not contain any announcement formulas and $S5$ is complete (see Theorem 1), we have $\vdash^{S5}t(\phi)$. Since PAL is an extension of S5, we also have $\vdash t(\phi)$ and so $\vdash\phi$. ∎
6 Examples of Public Announcements Revisited
We have established a B-eS for PAL, but it is important to see how this semantics can be used to analyze the connection between announcements and information. To this end, we discuss the two example puzzles, we have already given in Sections 3.1 and 3.2, respectively.
We have chosen these two examples because they highlight two different aspects of how information is treated differently in the inferentialist setup of base-extension semantics compared to the modal theoretic view of Kripke semantics.
Concretely, in the card game example we highlight how base rules can be used intensionally to construct bases that best correspond to the situation one is trying to model, in contrast to the extensional character of worlds and valuations in Kripke semantics.
In our discussion of the muddy children puzzle, we focus on investigating exactly what base rules are actually necessary to include in bases in order for the example to play out in the expected way or, phrased differently, consider what information do the children need in order to be able to find out that they are muddy.
In both cases, we give what we call minimal set-ups for a base-extension modelling of the examples. We use minimal here in the rather weak sense that any extension of the bases will result in the desired results. We do not make the stronger claim that these are the minimal possible requirements and any weaker set-up will fail. Although such a result would be desirable, it goes beyond the scope of this paper.
6.1 Three-player Card Game
As we have seen in Subsection 3.1, the Kripke semantics modeling for the card game example (Example 1) lines up exactly with our informal interpretation. There is, however, one aspect of the modeling we glossed over and that is the implementation of the rules of the game. Obviously, we want to show that we can model this example in our base-extension semantics and get the same intuitive results as the Kripke semantics. Given the equivalence between maximally-consistent bases and worlds in Kripke semantics, it should come as no surprise that simply taking maximally-consistent bases that correspond to the worlds used in the Kripke models works.
However, we can also give a more satisfying treatment of the rules of the card game, as base rules directly embedded into our bases. For this reason we want to construct our bases with as few base rules as possible resulting in a minimal model, in the sense given above. This will allow us to make the stronger claim that the necessary reasoning by the agents will work on any model for which these minimal restrictions hold.
As discussed above the basic rules of the card game are that a) every card is held by exactly one agent and b) every agent holds exactly one card. We start by constructing a base at which ‘at most’ versions of a) and b) hold. This base will be used as a subset for our final bases.
What we mean by ‘every card is held by at most one agent’ is that any base the players are considering at which a card is held by more than one agent must be inconsistent. So, for example, we want $(0_{a}\wedge 0_{b})\to\bot$ to hold at our base. Similarly, for ‘every agent holds at most one card’, we require $(0_{a}\wedge 1_{a})\to\bot$. So we want any combination of two agents holding the same cards and two cards being held by the same agent to lead to contradiction, that is to say cause every propositional atom to hold. We can straightforwardly create rule-schema that cover all those cases. We call the base for which this holds $\mathscr{B}_{cards}$ and give a formal definition as part of Proposition 1.
We cannot give similar base rules to establish the ‘at least’ parts of a) and b) directly. Given the form of our base rules, as given in Definition 10, it is not possible to give rules to that effect. However, similarly to the Kripke semantics version, we take bases at which the agents have each drawn a card and each card has been drawn.
There is one more ‘rule’ of the card game we have not yet discussed: Every player can only see their own hand. This is a modal rule that directly impacts what information the players can perceive. As such, this is implemented by our choice of $\mathfrak{R}_{A}$. Similarly to in Kripke semantics, we choose an $\mathfrak{R}_{a}$ at which Anne only considers bases at which they drew the same card. The same goes for Bob and Cath.
We do not claim that all rules for any game that can be modelled using PAL can be implemented either as a base rule or as a modal rule in constructing the relevant $\mathfrak{R}_{A}$, as we have seen with the ‘at most’ versions of (a) and (b). We do however, argue that in cases in which we can this provides a more faithful representation of the example as compared to the abstract choice of valuations in Kripke semantics.
Proposition 1.
For all $M,N,\in\{0,1,2\}$ with $M\neq N$, $i,j\in\{a,b,c\}$ with $i\neq j$, and $p$,
$\mathscr{B}_{cards}=\{(N_{i},N_{j}\rightarrow p),(N_{i},M_{i}\rightarrow p)\}$.
Let $\mathscr{B}_{MNP}$ be any consistent superset of $\mathscr{B}_{cards}$ s.t. $\Vdash_{\mathscr{B}_{MNP}}M_{a}\wedge N_{b}\wedge P_{c}$. For any set of modal relations $R_{A}$ with $A=\{a,b,c\}$ s.t. there are
$\mathscr{B}_{012},\mathscr{B}_{021},\mathscr{B}_{102},\mathscr{B}_{120},\mathscr{B}_{201}$ and $\mathscr{B}_{210}$
with
$\mathfrak{R}_{a}\mathscr{B}_{012}\mathscr{B}_{021}$, $\mathfrak{R}_{a}\mathscr{B}_{102}\mathscr{B}_{120}$, $\mathfrak{R}_{a}\mathscr{B}_{201}\mathscr{B}_{210}$,
$\mathfrak{R}_{b}\mathscr{B}_{102}\mathscr{B}_{201}$, $\mathfrak{R}_{b}\mathscr{B}_{012}\mathscr{B}_{210}$, $\mathfrak{R}_{b}\mathscr{B}_{021}\mathscr{B}_{120}$,
$\mathfrak{R}_{c}\mathscr{B}_{120}\mathscr{B}_{210}$, $\mathfrak{R}_{c}\mathscr{B}_{021}\mathscr{B}_{201}$, $\mathfrak{R}_{c}\mathscr{B}_{012}\mathscr{B}_{102}$,
and for which these, together with the necessary relations to make these $S5$ (i.e., reflexive, transitive, and Euclidean), are the only relations involving these bases,
$\Vdash_{\mathscr{B}_{012},\mathfrak{R}_{A}}[\neg 1_{a}]K_{c}(0_{a}\wedge 1_{b}\wedge 2_{c})$.
Note that besides building the subset of the bases we require to model the card game example, $\mathscr{B}_{cards}$ also represents the state of the game before any card has been drawn.
By Definition 16, $\Vdash_{\mathscr{B}_{012},\mathfrak{R}_{A}}[\neg 1_{a}]K_{c}(0_{a}\wedge 1_{b}\wedge 2_{c})$ iff for all $\mathscr{C}_{012}\supseteq\mathscr{B}_{012}$, if $\Vdash_{\mathscr{C}_{012},\mathfrak{R}_{A}}\neg 1_{a}$, then $\Vdash_{\mathscr{B}_{012},\mathfrak{R}_{A}}^{\neg 1_{a}}K_{c}(0_{a}\wedge 1_{b}\wedge 2_{c})$. As $\mathfrak{R}_{c}$ is a modal relation we know that any $\mathscr{D}$, s.t. $\mathfrak{R}_{c}\mathscr{C}_{012}\mathscr{D}$, has to be a superset of $\mathscr{B}_{012}$ or $\mathscr{B}_{102}$, by condition (d) of Definition 11. We call $\mathscr{D}_{012}$ those $\mathscr{D}\supseteq\mathscr{B}_{012}$ and $\mathscr{D}_{102}$ those $\mathscr{D}\supseteq\mathscr{B}_{102}$. We cannot have any $\mathfrak{R}^{\prime}_{A}\in\mathfrak{R}_{A}|1_{a}^{\mathscr{C}_{012}}$ with $\mathfrak{R}^{\prime}_{c}\mathscr{C}_{012}\mathscr{D}_{102}$, as $\nVdash_{\mathscr{D}_{102},\mathfrak{R}_{A}}\neg 1_{a}$. So we only need to consider $\mathscr{D}_{012}$. As $\mathscr{D}_{012}\supseteq\mathscr{B}_{012}$, we have $\Vdash_{\mathscr{D}_{012},\mathfrak{R}_{A}}^{1_{a}}0_{a}\wedge 1_{b}\wedge 2_{c}$ and so also $\vdash_{\mathscr{B}_{012},\mathfrak{R}_{A}}[\neg 1_{a}]K_{c}(0_{a}\wedge 1_{b}\wedge 2_{c})$. So the example plays out in exactly the same way as we expected it to.
In fact the basic structure of the model remains largely unchanged from the Kripke semantics version illustrated in Figure 1. Only now we have to consider supersets and sometimes multiple such supersets for the same minimal base (e.g., $\mathscr{B}_{012}$). However, due to condition (d) of modal relations, we know that no other bases are reachable by $\mathfrak{R}_{A}$ and so the example plays out in the same way. We illustrate this in Figure 12.
$\{102\}$$\{201\}$$\mathscr{C}_{012}+\{012\}$$\{120\}$$\{210\}$$\{021\}$$a$$a$$a$$b$$b$$b$$c$$c$$c$$\{201\}$$\mathscr{C}_{012}+\{012\}$$\{210\}$$\{021\}$$a$$a$$b$$c$
Figure 12: The structure of $\mathfrak{R}_{A}^{\mathscr{C}_{012}}$ (left) and any $\mathfrak{R}^{\prime}_{A}\in\mathfrak{R}_{A}|\neg 1_{a}^{\mathscr{C}_{012}}$ (right) for bases $\mathscr{C}_{012}\supseteq\mathscr{B}_{012}$ with $\{012\}$ standing for supersets of $\mathscr{B}_{012}$ etc.
6.2 The Muddy Children Puzzle
We model the muddy children puzzle (Example 2) for three children in which two, Anne and Bob, are muddy in the base-extension semantics for PAL. As with the card game example, we want to give a minimal example. Unlike the card game, however, we do not have any underlying rules of the example that can be implemented into base rules of our example. We simply have one base for each of the possible combinations of muddy children. The fact that the children cannot see their own forehead is, like the players only being able to see their own card, expressed by the corresponding model relations between those bases.
We use the idea of using minimal bases to once again show the necessary information to get the wanted result. The question we want to investigate is exactly what base rules are necessary for the muddy children to be able to draw the expected conclusions about which of them are, in fact, muddy. For bases in which a specific child, say Anne, is muddy it turns out that it is not enough to for ‘Anne is muddy’ to not be supported, but rather we require for Anne to not be muddy (i.e., the negation has to be supported). We show this by showing that Proposition 2, in which we do not make this requirement, fails.
Proposition 2.
For a set of agents $A=\{a,b,c\}$ and any $N\in P(A)$ (i.e., the powerset of $A$), let $\mathscr{B}_{N}$ be any consistent base such that for all $j\in A$ with $m_{j}\in N$, $\Vdash_{\mathscr{B}_{N},\mathfrak{R}_{A}}m_{j}$.
For any set of modal relations $R_{A}$ with $A=\{a,b,c\}$ s.t. there are
,$\mathscr{B}_{\emptyset},\mathscr{B}_{\{a\}},\mathscr{B}_{\{b\}},\mathscr{B}_{\{c\}},\mathscr{B}_{\{a,b\}},\mathscr{B}_{\{a,c\}},\mathscr{B}_{\{b,c\}}$ and $\mathscr{B}_{\{a,b,c\}}$ with
$\mathfrak{R}_{a}\mathscr{B}_{\emptyset}\mathscr{B}_{\{a\}}$, $\mathfrak{R}_{a}\mathscr{B}_{\{b\}}\mathscr{B}_{\{a,b\}}$, $\mathfrak{R}_{a}\mathscr{B}_{\{c\}}\mathscr{B}_{\{a,c\}}$, $\mathfrak{R}_{a}\mathscr{B}_{\{b,c\}}\mathscr{B}_{\{a,b,c\}}$,
$\mathfrak{R}_{b}\mathscr{B}_{\emptyset}\mathscr{B}_{\{b\}}$, $\mathfrak{R}_{b}\mathscr{B}_{\{a\}}\mathscr{B}_{\{a,b\}}$, $\mathfrak{R}_{b}\mathscr{B}_{\{c\}}\mathscr{B}_{\{b,c\}}$, $\mathfrak{R}_{b}\mathscr{B}_{\{a,c\}}\mathscr{B}_{\{a,b,c\}}$,
$\mathfrak{R}_{c}\mathscr{B}_{\emptyset}\mathscr{B}_{\{c\}}$, $\mathfrak{R}_{c}\mathscr{B}_{\{a\}}\mathscr{B}_{\{a,c\}}$, $\mathfrak{R}_{c}\mathscr{B}_{\{b\}}\mathscr{B}_{\{b,c\}}$, $\mathfrak{R}_{c}\mathscr{B}_{\{a,b\}}\mathscr{B}_{\{a,b,c\}}$,
and for which those, together with the necessary relations to make these $S5$ (i.e., reflexive, transitive, and Euclidean), are the only relations involving these bases, we have
$\Vdash_{\mathscr{B}_{\{a,b\}},\mathfrak{R}_{A}}[(m_{a}\vee m_{b}\vee m_{c})][\neg(K_{a}m_{a}\vee K_{a}\neg m_{a})\wedge\neg(K_{b}m_{b}\vee K_{b}\neg m_{b})\wedge\neg(K_{c}m_{c}\vee K_{c}\neg m_{c})](K_{a}m_{a}\wedge K_{b}m_{b})$.
To show that Proposition 2 fails, we give a counterexample. To make things simpler and avoid any considerations of the sub- and superset relations, for every $\mathscr{B}_{N}$, let $p_{N}$ be a propositional atom s.t. the base rule $(\Rightarrow p_{N})\in\mathscr{B}_{N}$ and, for all $\mathscr{B}_{M}$ s.t. $M\neq N$, $(\Rightarrow p_{N})\notin\mathscr{B}_{M}$. To further keep the example as simple as possible, we add simply the rules corresponding to $(\Rightarrow m_{i})$ for all $i\in N$ to a base $\mathscr{B}_{N}$. So, for example, $\mathscr{B}_{\{a,b\}}=\{(\Rightarrow p_{\{a,b\}}),(\Rightarrow m_{a}),(\Rightarrow m_{b})\}$. We also still let $\phi=(m_{a}\vee m_{b}\vee m_{c})$ and $\psi=\neg(K_{a}m_{a}\vee K_{a}\neg m_{a})\wedge\neg(K_{b}m_{b}\vee K_{b}\neg m_{b})\wedge\neg(K_{c}m_{c}\vee K_{c}\neg m_{c})$.
For $\mathscr{B}_{\emptyset}$, we are taking something a bit different. Remember that the conditions above are minimal conditions and so we can add base rules to our bases unless they contradict them directly. So let =$\mathscr{B}_{\emptyset}=\{(\Rightarrow p_{\emptyset}),(p_{\emptyset}\Rightarrow m_{a})\}$.
Given these bases it is easy to show that the example will not play out in the way we expect it to. Note that now $\Vdash_{\mathscr{B}_{\emptyset},\mathfrak{R}_{A}}m_{a}$ and so the father’s initial announcement of $\phi=m_{a}\vee m_{b}\vee m_{c}$ that at least one child is muddy fails to eliminate the base $\mathscr{B}_{\emptyset}$ (see Figure 13).
$\mathscr{B}_{\emptyset}$$\mathscr{B}_{\{b\}}$$\mathscr{B}_{\{c\}}$$\mathscr{B}_{\{a\}}$$\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $a$$b$$c$$b$$c$$a$$c$$a$$b$$c$$b$$a$Longrightarrow $\mathscr{B}_{\emptyset}$$\mathscr{B}_{\{b\}}$$\mathscr{B}_{\{c\}}$$\mathscr{B}_{\{a\}}$$\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $a$$b$$c$$b$$c$$a$$c$$a$$b$$c$$b$$a$
Figure 13: $S^{*}_{A}$ on the left and $S^{*}_{A}|\phi^{\mathscr{B}_{\{a,b\}}}$ (Definition 18) on the right. The announcement of $\phi$ fails to eliminate $\mathscr{B}_{\emptyset}$ and the relations to it (highlighted by dashed lines).
It should already be clear that the modelling of the example will fail, however, it is useful to see how exactly this failure is going to impact the result.
Normally, we should expect the announcement of $\psi$ to get rid of the bases $\mathscr{B}_{\{a\}},\mathscr{B}_{\{b\}},$ and $\mathscr{B}_{\{c\}}$, because if at least one child is muddy, in these bases the sole muddy child would realize that it is muddy. How does that change with $\mathscr{B}_{\emptyset}$ still around? Note that at the bases $\mathscr{B}_{\emptyset}$ and $\mathscr{B}_{\{a\}}$, Anne only has access to bases in which $m_{a}$ is supported. By Lemma 8, we know that this will also hold for all superset bases. So ,$\Vdash^{\phi}_{\mathscr{B}_{\emptyset},\mathfrak{R}_{A}}K_{a}m_{a}$ and $\Vdash^{\phi}_{\mathscr{B}_{\{a\}},\mathfrak{R}_{A}}K_{a}m_{a}$. This means that $\psi$ fails at these bases. At all other bases, however, the children do not yet know whether they are muddy and so $\psi$ continues to hold. So, the announcement of $\psi$ eliminates $\mathscr{B}_{\emptyset}$ and $\mathscr{B}_{\{a\}}$ and leaves us with the model in Figure 14.
$\mathscr{B}_{\emptyset}$$\mathscr{B}_{\{b\}}$$\mathscr{B}_{\{c\}}$$\mathscr{B}_{\{a\}}$$\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $a$$b$$c$$b$$c$$a$$c$$a$$b$$c$$b$$a$Longrightarrow $\mathscr{B}_{\{b\}}$$\mathscr{B}_{\{c\}}$$\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $a$$c$$a$$b$$c$$b$$a$
Figure 14: After the announcement of $\psi$ (right side), the bases $\mathscr{B}_{\{b\}}$ and $\mathscr{B}_{\{c\}}$ fail to be eliminated (with the corresponding relations highlighted by dashed lines).
Remember that we wanted to show that $\Vdash_{\mathscr{B}_{\{a,b\}},\mathfrak{R}_{A}}[\phi][\psi](K_{a}m_{a}\wedge K_{b}m_{b})$. As it turns out $\Vdash^{\phi,\psi}_{\mathscr{B}_{\{a,b\}},\mathfrak{R}_{A}}K_{b}m_{b}$, but $\nVdash^{\phi,\psi}_{\mathscr{B}_{\{a,b\}},\mathfrak{R}_{A}}K_{a}m_{a}$.
Interestingly, Anne being muddy at base $\mathscr{B}_{\emptyset}$ did not change Bob’s ability to find out whether they are muddy or not. Similarly, Cath’s reasoning remains largely unchanged in our example as well, even though they were unable to conclude whether they are muddy or not. After the announcements of $\phi$ and $\psi$, Cath still considers $\mathscr{B}_{\{a,b\}}$ and $\mathscr{B}_{\{a,b,c\}}$. This suggests that it is important for Anne that we require $\mathscr{B}_{\emptyset}$ to not support $m_{a}$.
Given that we are trying to establish minimal conditions for these bases, there is an important difference between not having ‘Anne is muddy’ and ‘Anne is not muddy’. In order for the announcement of $\phi$ to have its intended effect $\mathscr{B}_{\emptyset}$ has to be a base at which none of the children is muddy. Similarly, for all the bases that represent cases in which a child is not muddy it is not enough to not require the child to be muddy. Take a base $\mathscr{B}_{\{b\}}$ at which $m_{a}$ is supported. At the bases $\mathscr{B}_{\{b\}}$ and $\mathscr{B}_{\{a,b\}}$, Anne now only has access to bases at which they are muddy and so $K_{a}m_{a}$ will be supported even before an announcement even takes place and $\psi$ cannot be announced at $\mathscr{B}_{\{a,b\}}$ [^3] Even though this does not technically cause Proposition 2 to fail, as $K_{a}m_{a}\wedge K_{b}m_{b}$ will vacuously hold when $\psi$ cannot be announced, it shows that the model does not behave as expected. Consider that $K_{c}m_{c}$ will also hold even though Cath is not muddy..
So, in order to give actual minimal conditions for bases corresponding to the muddy children puzzle, we update Proposition 2 by requiring the children to be not muddy at the relevant bases.
Proposition 3.
For a set of agents $A=\{a,b,c\}$ and any $N\in P(A)$ (i.e., the powerset of $A$), let $\mathscr{B}_{N}$ be any consistent base such that for all $j\in A$ with $m_{j}\in N$, $\Vdash_{\mathscr{B}_{N},\mathfrak{R}_{A}}m_{j}$ and for all $j\in A$ s.t. $m_{j}\notin N$, $\Vdash_{\mathscr{B}_{N},\mathfrak{R}_{A}}m_{j}\to\bot$. For any set of modal relations $R_{A}$ with $A=\{a,b,c\}$ s.t. there are
,$\mathscr{B}_{\emptyset},\mathscr{B}_{\{a\}},\mathscr{B}_{\{b\}},\mathscr{B}_{\{c\}},\mathscr{B}_{\{a,b\}},\mathscr{B}_{\{a,c\}},\mathscr{B}_{\{b,c\}}$ and $\mathscr{B}_{\{a,b,c\}}$
with
$\mathfrak{R}_{a}\mathscr{B}_{\emptyset}\mathscr{B}_{\{a\}}$, $\mathfrak{R}_{a}\mathscr{B}_{\{b\}}\mathscr{B}_{\{a,b\}}$, $\mathfrak{R}_{a}\mathscr{B}_{\{c\}}\mathscr{B}_{\{a,c\}}$, $\mathfrak{R}_{a}\mathscr{B}_{\{b,c\}}\mathscr{B}_{\{a,b,c\}}$,
$\mathfrak{R}_{b}\mathscr{B}_{\emptyset}\mathscr{B}_{\{b\}}$, $\mathfrak{R}_{b}\mathscr{B}_{\{a\}}\mathscr{B}_{\{a,b\}}$, $\mathfrak{R}_{b}\mathscr{B}_{\{c\}}\mathscr{B}_{\{b,c\}}$, $\mathfrak{R}_{b}\mathscr{B}_{\{a,c\}}\mathscr{B}_{\{a,b,c\}}$,
$\mathfrak{R}_{c}\mathscr{B}_{\emptyset}\mathscr{B}_{\{c\}}$, $\mathfrak{R}_{c}\mathscr{B}_{\{a\}}\mathscr{B}_{\{a,c\}}$, $\mathfrak{R}_{c}\mathscr{B}_{\{b\}}\mathscr{B}_{\{b,c\}}$, $\mathfrak{R}_{c}\mathscr{B}_{\{a,b\}}\mathscr{B}_{\{a,b,c\}}$,
and for which these, together with the necessary relations to make these $S5$ (i.e., reflexive, transitive, and Euclidean), are the only relations involving these bases,
$\Vdash_{\mathscr{B}_{\{a,b\}},\mathfrak{R}_{A}}[(m_{a}\vee m_{b}\vee m_{c})][\neg(K_{a}m_{a}\vee K_{a}\neg m_{a})\wedge\neg(K_{b}m_{b}\vee K_{b}\neg m_{b})\wedge\neg(K_{c}m_{c}\vee K_{c}\neg m_{c})](K_{a}m_{a}\wedge K_{b}m_{b})$.
It is straightforward to see that for any $\mathscr{B}_{\emptyset}$, we have $\nVdash_{\mathscr{B}_{\emptyset},R_{A}}\phi$ and so it is eliminated in the way we expect by the announcement of $\phi$ (Figure 15).
$\mathscr{B}_{\emptyset}$$\mathscr{B}_{\{b\}}$$\mathscr{B}_{\{c\}}$$\mathscr{B}_{\{a\}}$$\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $a$$b$$c$$b$$c$$a$$c$$a$$b$$c$$b$$a$Longrightarrow $\mathscr{B}_{\{b\}}$$\mathscr{B}_{\{c\}}$$\mathscr{B}_{\{a\}}$$\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $b$$c$$a$$c$$a$$b$$c$$b$$a$
Figure 15: $S^{*}_{A}$ on the left and $S^{*}_{A}|\phi^{\mathscr{B}_{\{a,b\}}}$ on the right (Definition 18). Given the new requirement in Proposition 3, the announcement of $\phi$ goes as expected.
Similarly, $\psi$ will fail at the bases $\mathscr{B}_{\{a\}}$, $\mathscr{B}_{\{b\}}$, and $\mathscr{B}_{\{c\}}$ after the announcement of $\phi$ as the child that is muddy at them will know that they are muddy. This leaves us with the final result in Figure 16. Importantly (and as expected), $\Vdash_{\mathscr{B}_{\{a,b\}},\mathfrak{R}_{A}}[\phi][\psi](K_{a}m_{a}\wedge K_{b}m_{b})$.
$\mathscr{B}_{\{b\}}$$\mathscr{B}_{\{c\}}$$\mathscr{B}_{\{a\}}$$\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $b$$c$$a$$c$$a$$b$$c$$b$$a$Longrightarrow $\mathscr{B}_{\{a,b\}}$$\mathscr{B}_{\{a,c\}}$$\mathscr{B}_{\{b,c\}}$$\mathscr{B}_{\{a,b,c\}}$ $c$$b$$a$
Figure 16: $S^{*}_{A}|\phi^{\mathscr{B}_{\{a,b\}}}$ on the left and $S^{*}_{A}|(\phi,\psi)^{\mathscr{B}_{\{a,b\}}}$ on the right (Definition 18). The announcement of $\psi$ also functions as intended.
It should be reiterated here that this only proves that the conditions in Proposition 3 guarantee that the muddy children puzzle goes through in our P-tS-based setting. It does not prove that these are in fact minimal requirements in the strong sense. There can be other requirements, either weaker or simply different, that will also have the desired results. We do contend, however, that these conditions are at least the most natural restrictions to put on the bases and directly represent the situation as given in the description of the example.
We suggest that the need for imposing such conditions in the P-tS-based analysis of the muddy children puzzle is wholly consistent with an inferentialist position in which we expect to make assumptions explicit (cf. the position presented in [2]).
6.3 Information and Intensionality
When trying to analyse information through examples like the card game and the muddy children, it is important to identify exactly what it is that allows the agents to draw their conclusions. This means that in the context of public announcements we must identify exactly what about the announcement and the context in which it was given/received allows the agents to make their specific inferences.
In Section 3, we can see that the way in which worlds in the Kripke semantics modelling each of these examples have been chosen is purely extensional. For the card game, we simply picked worlds at which the card-draw lined up with the example. Similarly, in the muddy children example, we picked a world that corresponds to each configuration of the muddy children for that example. In fact, worlds are inherently extensional in this way because, ultimately, the truth of a formula is based on such abstract interpretations.
Our inferentialist analysis aims to achieve a deeper understanding of the information involved in these examples is characterized by drawing on the intensional character of bases. As we have already mentioned at the beginning of this section, the two examples highlight two ways in which bases and base rules can be used. After going through the examples, we want to restate these two ways and connect them with an intensional approach to information.
In the card game example, we specifically do not just consider bases at which the rules of the game hold, but rather construct bases that have base rules which directly represent those rules of the game. This allows us to be able to say not just that these rules hold, as is the case in the Kripke model, but exactly why they hold.
In both examples, we give a minimal set-up that tells us that any extension of this set-up will allow the example to go through as expected. This is especially highlighted in the muddy children example, in which the example fails for a specific set of requirements. This allows us to identify exactly what bases and relations are required for the example to go through and so what information must be encoded in the model for the agent to make the expected inferences. Although maybe not as obviously intensional in character as implementing the rules directly, as is done in the card game, any minimal instantiation of this set-up by specific base rules will say a child is muddy (or not) is supported exactly because of those base rules. Furthermore, at any extension of such a minimal instantiation, that support persists because of the presence of those same base rules. [^4] Of course, for any extension that is also an extension of another minimal set-up, there are now two reasons for the formula to be supported. In this sense, smaller bases are ‘more’ intensional than larger ones. At the extreme, maximally-consistent bases, which can be considered to amount to Kripke worlds, contain all the ways that cause the formula to be supported and, therefore, are extensional in character.
7 Future Work
We view this work as a first step in an inferentialist analysis of the group of dynamic epistemic logics. Public announcement logic is the simplest of these logics. The natural next step is to look at action model logics (AML). In our setting, public announcement leave the bases unchanged and only change the modal relations. This means that a base before an action is the same as after an action. In AML, there can be multiple possible actions that the agents consider could have taken place. So, a base before an action would have to correspond to multiple bases after the action. There is no straightforward way of generalizing our semantics to incorporate that. We suggest an approach that instead allows the agents to consider different modal relations, one for each possible action they consider. In this sense the agents uncertainty about what action actually took place is modelled by their uncertainty in what are the correct modal relations.
There are also some interesting aspects of (static) epistemic logic and PAL that still need to be analysed from the inferentialist perspective. We have yet to discuss group notions of knowledge, such as distributed knowledge or common knowledge. Common knowledge is especially interesting in the context of PAL, as there is a close association between it and public announcements. Another notion relevant to PAL is unsuccessful update; that is, announcements of formulae that become false when they are announced. All this analysis is important in identifying the informational content of announcements. See [4] for an overview on these concepts in the context of dynamic epistemc logic in general.
We suggest that the future work discussed in this section is instrumental in providing an inferentialist analysis of information in certain circumstances. Specifically, we conjecture that the inferentialist view of meaning as being derived from use — in particular, from inference in P-tS — provides a fruitful perspective on the concept of information itself, not only from a foundational perspective, but also as it occurs in settings such as informatics and technology. While this work and the directions for future work discussed above are useful in their own terms, they also provide some first steps towards an inferentialist account of information.
Acknowledgements
This work has been partially supported by the UK EPSRC grants EP/S013008/1 and EP/R006865/1 and by the EU MOSAIC MCSA-RISE project. We are grateful to Gabriele Brancati, Tim Button, Yll Buzoku, Alex Gheorghiu, Tao Gu, Timo Lang, Elaine Pimentel, and Will Stafford for helpful discussions of this work.